### v1

# Which Similarity-Sensitive Entropy?

Abstract
A canonical step in quantifying a system is to measure its entropy. Shannon entropy and other traditional entropy measures capture only the information encoded in the frequencies of a system’s elements. Recently, Leinster, Cobbold, and Reeve (LCR) introduced a method that also captures the rich information encoded in the similarities and differences among elements, yielding similarity-sensitive entropy. More recently, the Vendi score (VS) was introduced as an alternative, raising the question of how LCR and VS compare, and which is preferable. Here we address these questions conceptually, analytically, and experimentally, using 53 machine-learning datasets. We show that LCR and VS can differ by orders of magnitude and can capture complementary information about a system, except in limiting cases. We demonstrate that both LCR and VS depend on how similarities are scaled and introduce the concept of “half distance” to parameterize this dependence. We prove that VS provides an upper bound on LCR for several values of the Rényi-Hill order parameter and conjecture that this bound holds for all values. We conclude that VS is preferable only when interpreting elements as linear combinations of a more fundamental set of “ur-elements” or when the system or dataset possesses a quantum-mechanical character. In the broader circumstance where one seeks simply to capture the rich information encoded by similarity, LCR is favored; nevertheless, for certain half-distances the two methods can complement each other.
1 Introduction
Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across science, engineering, commerce, and law [leinster2016maximizing, Shannon1951, Montemurro1994, tawilCompetitionCaliforniasMediCal2022]. Traditional entropy, as formulated by Shannon [shannon1948mathematical] and generalized by Rényi [renyiMeasuresEntropyInformation1961] (and later Tsallis [tsallis1988possible]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits or nats required to encode the shape of the distribution, with smaller values for more skew distributions and larger values for flatter ones (Fig. 1). These measures rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig. 1).
Figure 1: The concepts of element, frequency distribution, traditional entropy (at $q=1$), similarity function, similarity measure ($z_{i}j$, with examples), similarity matrix ($Z$), and similarity-sensitive entropy (here, LCR, also at $q=1$). Entropy values are expressed in effective-number form, i.e. in units of effective number of images present in the dataset. Different similarity measures can be chosen (see Section 5.2); in this illustrative example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
Incorporating similarities into the calculation of entropy results in so-called similarity-sensitive entropy. This was first accomplished by Leinster and Cobbold [leinster2012measuring] and extended by Reeve and colleagues [reeve2014partition], in a framework we refer to here as LCR [leinsterEntropyDiversityAxiomatic2020]. By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig. 1). This capability is valuable in many domains [nguyen2023greylock], especially in machine learning (ML) [couch2024beyond, couch2025x] (for example, to achieve state-of-the-art performance more efficiently [chinnENRICHingMedicalImaging2023]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e. similarity-insensitive) entropies because their frequency distributions are both flat. Similarity-sensitive entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [arnaoutFutureBloodTesting2021, aroraRepertoirescaleMeasuresAntigen2022], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat and therefore traditional entropies less informative.
More recently, an alternative form of similarity-sensitive entropy has been described, consisting of the Vendi score [friedman2022vendi, pasarkar2025vendiscope] and its variants [pasarkar2023cousins]—hereafter collectively VS—separate from LCR. This raises the practical question of which form of similarity-sensitive entropy to choose for a given situation: LCR or VS. The current work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets used for ML, and comparing their values.
2 Related Work
Entropies fall into two classes: traditional/similarity‑insensitive and similarity‑sensitive. Although one often sees reference to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), entropies are actually families of measures; individual members are distinguished by how they weight frequencies and, for similarity-sensitive entropy, by their similarity matrix (see Table 1).
2.1 Traditional entropy (similarity-insensitive)
[…]
Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the system [jost2006entropy]. Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types, while larger $q$ give greater weight to common types. $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form. Hill demonstrated that several other familiar statistics are special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [hill1973diversity]. For this reason, Hill’s formulation is often described as a unifying framework. $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
$D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1.\end{cases}$ (1)
where $\mathbf{p}$ denotes the frequency distribution of unique elements. Other generalizations, such as Tsallis entropy [tsallis1988possible], also exist. What these traditional entropies all have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
2.2 Similarity-sensitive entropy
2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
The LCR framework [leinsterEntropyDiversityAxiomatic2020] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropy does not. A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$. The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1. The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive Rényi entropies $H_{\alpha}^{Z}$ and are given by
$D_{q}^{Z}(\mathbf{p};Z)=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}$ (2)
Here $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency‑weighted average similarity of element $i$ to all the elements (including itself). When this average is large, element $i$ is considered ordinary, making Eq. 2 interpretable as the average “ordinariness” [leinster2012measuring] across all elements. $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
  • phylogenetic diversity [chao2010phylogenetic] in the special case where $z_{ij}$ forms an ultrametric, typically derived from a phylogenetic tree;
  • functional diversity [chaoAttributediversityApproachFunctional2019] when the similarity pertains to elements’ function, for example the binding similarity between pairs of antibodies or TCRs [aroraRepertoirescaleMeasuresAntigen2022];
  • attribute diversity [chaoAttributediversityApproachFunctional2019] when interpreting the system as a set of attribute contributions instead of frequencies; and
  • similarity-sensitive or similarity-aware diversity [nguyen2023greylock] more generally.
LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become uninformative. Representative applications in the life sciences include high‑throughput immunology (immunomes) [aroraRepertoirescaleMeasuresAntigen2022], microbiome research (metagenomics) [nguyen2023greylock], and medical imaging [couch2024beyond]. specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g. images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [couch2024beyond].
2.2.2 The Vendi score and its “cousins” (VS)
VS constitutes a related but separate class of similarity‑sensitive entropy measures. Like LCR, VS entropies are functions of a similarity matrix, but the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element. We can define $Z_{n}$ in terms of $Z$ as
$\left(Z_{n}\right)_{i,j}=Z_{s(i),s(j)}.$ (3)
where $s(i)$ is the unique element of which the $i^{\mathbf{th}}$ overall element is an example.
The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\lambda_{i}$ of $Z_{n}/n$. The division by $n$ normalizes $Z_{n}$ to unit trace (because as in $Z$, the diagonals of $Z_{n}$ equal 1):
$\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\right)$ (4)
VS is to the similarity matrix what the von Neumann entropy is to the quantum density matrix. Replacing the Shannon term in Eq. 4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [pasarkar2023cousins] of the Vendi score:
$\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\lambda_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
$\mathrm{VS}_{q}$ is a 1-parameter family of scores with $q$ as in $D_{q}^{Z}$.
Note that LCR can also be written as a function of $Z_{n}$ if desired:
$D_{q}^{Z}(\mathbf{p};Z)=D_{q}^{Z}(\tilde{\mathbf{p}}=\frac{1}{n};Z_{n})$ (6)
where $\tilde{\mathbf{p}}$ is the uniform distribution on the $n$ elements.
Entropy type and formula Frequency weighting Similarity-sensitive? Notes and applications
Shannon (Boltzmann–Gibbs)
$H_{1}(\mathbf{p})=-\sum_{i=1}^{S}p_{i}\log p_{i}$
$q=1$ No Information theory, Ecology, ML loss functions, Thermodynamics
Rényi entropy ($\alpha$)
$H_{\alpha}(mathbf{p})=\frac{1}{1-\alpha}\log\sum_{i}p_{i}^{\alpha}$
Any $q$ No Info-theoretic security, Ecology, Fractal analysis, Physics
Tsallis entropy ($q$-entropy)
$S_{q}(\mathbf{p})=\frac{1}{q-1}\big(1-\sum_{i}p_{i}^{q}\big)$
Any $q$ No Non-extensive statistical mechanics, Turbulence, Astrophysics
Quantum (von Neumann) entropy
$S(\rho)=-\operatorname{Tr}(\rho\log\rho)$
$q=1$ No Quantum information, Entanglement, Quantum thermodynamics
LCR $D_{q}^{Z}(\mathbf{p};Z)=$
$\big(\sum_{i}p_{i}(Z\mathbf{p})_{i}^{q-1}\big)^{1/(1-q)}$ for $q\neq 1$
$\exp(-\sum_{i}p_{i}\log(Z\mathbf{p})_{i})$ for $q=1$
Any $q$ Yes Phylogenetic, attribute, and functional diversity are special cases. Biodiversity, microbiomes, language…
Vendi score
[…]
$q=1$ Yes $\lambda_{i}$ are eigenvalues of $Z_{n}=Z/N$ (positive semi-definite). ML diversity, Generative model evaluation, Ecology
Cousins of the Vendi score
$\text{VS}_{q}(\mathbf{x},\mathbf{k})=\left(\sum_{i}\lambda_{i}^{q}\right)^{\tfrac{1}{1-q}}$
Any $q$ Yes Quantum-inspired stats, Class diversity profiling
Other variants
Burg: $\sum_{i}\log p_{i}$
KL: $D_{KL}(p\|q)=\sum_{i}p_{i}\log\frac{p_{i}}{q_{i}}$
CRE: $\int_{0}^{\infty}\bar{F}(x)\log\bar{F}(x)\,dx$
Partly Partly Spectral estimation (Burg); Bayesian inference (KL); Survival analysis (CRE)
Table 1: Summary of major entropy families, their frequency-weighting forms, similarity sensitivity, and main applications.
3 Methods
3.1 Datasets and definitions
The following imaging and tabular datasets were downloaded for benchmarking LCR and VS:
  • Image datasets. 22 large ($\geq 10,000$-image) standard vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [yangMedMNISTClassificationDecathlon2021a]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [visipediaInat_comp2017READMEmd]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [mei2022radimagenet], the NIH ChestXRay dataset [wangChestXRay8HospitalScaleChest2017]; COCO (Common Objects in Context) [linMicrosoftCOCOCommon2015]; and MIDRC COVIDx CXR-4 [wuCOVIDxCXR4Expanded2023]. For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
  • Tabular datasets. After selecting for a maximum of 30 columns and 200,000 rows, the 31 most popular datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [lucieImprovedPython2024]. For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
Definitions. The terms “system” and “dataset” are used interchangeably. The elements of image and tabular datasets are single images and single rows, respectively. Each element appears only once in these datasets; i.e. each element is also a unique element.
3.2 Similarity matrices
Similarity between each two (unique) elements $\textbf{x}_{i}$ and $\textbf{x}_{j}$ was calculated from the Euclidean distance (L2 norm) as follows:
[…]
where $k$ is a parameter that was varied during experiments to test for effects on results. This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [leinsterEntropyDiversityAxiomatic2020]. Imaging datasets were first embedded into a two-dimensional space using UMAP [mcinnesUMAPUniformManifold2020] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq. 7 equivalent to the root-mean-square distance (RMSD).
3.3 Entropy calculations
The greylock Python package [nguyen2023greylock] was used to compute LCR as the $\gamma$ diversity; this is the effective number form of the corresponding entropy (see Section 6). The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with VS. VS was calculated using the vendi‑score package [vendi-score]. This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR. All computations were performed on a workstation equipped with multi‑core CPUs (greylock leverages parallelization).
3.4 Clustering
HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [mcinnesHdbscanHierarchicalDensity2017] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset. HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
3.5 Numerical experiments
Following a previous approach [Couch:2023pav], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize. A negative value indicates a counterexample. To meet the assumptions of this conjecture, we constructed $Z$ as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$. To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative. We assumed a uniform distribution of unique elements. Thus, we parameterized $Z$ by a $r\times n$ matrix $x$, from which we defined
$Z=w^{\intercal}w;w_{ij}=e^{x_{ij}}(\displaystyle\sum_{k}e^{2x_{kj}})^{-1/2}$ (8)
We minimized the difference $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for each combination of $q\in\{-10$, 2.7-2.7, 2-2, 1-1, 0.3-0.3, 0, $0.3$, 11, $2.7$, $3.4$, 44, 55, $2\pi$, $7$, 88, $9.1$, 10}10\}, $n\in\{3,5,10,20\}$, $r\in\{2,\lfloor n/2\rfloor,n\}$ if $n>3$ and $r\in\{2,3\}$ for $n=3$ (to avoid $r=1$). We also vary the distribution from which the initial guess fed to scipy.optimize was drawn. In total, 4420 minimization were done. The full Python code for performing these minimizations can be found in section A.1
4 Results
4.1 LCR and VS of major imaging and tabular ML datasets
We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits; the results below are for a random 10,000-image subset, taken for computational convenience. Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
Fig. 2a displays a UMAP embedding of the data, coloring each digit by its label. HDBSCAN detects six clusters, grouping together digits that appear visually similar (Fig. 2b). LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct (Fig. 2c). Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster. Taken together, these considerations suggest that LCR will approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose. Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10). Mathematically, this indicates that this MNIST subset possesses the same entropy as a hypothetical dataset containing 12-13 completely dissimilar elements (i.e. with zero pairwise similarity). This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the similarity-sensitive entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease the entropy relative to a scenario in which all 10 digit classes were entirely distinct.
In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements. However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑images (or “ur‑images”), each actual image being a linear combination of these ur-images. Figure 3 displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy. Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits. VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues. The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small.
Figure 2: The first 10,000 images of the MNIST digits dataset colored (a) by digit (labeled using representative images), (b) by HDBSCAN cluster, and (c) by similarity to one of the images (a “9”). The effective number of images in this dataset is 12.5 by LCR and 95.9 by VS.
Figure 3: (a) Top 100 eigenimages for the MNIST digits dataset and (b) their eigenvalues (in the gray region), alongside the full eigenvalue spectrum. Note the log axes.
Figure 4: LCR and VS for (a) imaging and (b) tabular datasets, sorted by VS ($q=1$; default $k$s).
Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets.
4.2 Trends with LCR and VS
Fig. 5 presents comparisons among LCR, VS, and the number of HDBSCAN clusters across all datasets, with imaging and tabular datasets evaluated separately. As observed for MNIST, both LCR and VS were generally larger than the cluster count. The principal exceptions were the CIFAR image datasets, where the absence of meaningful substructure produced hundreds of very small clusters. Correlations between the number of clusters and LCR, and between the number of clusters and VS, were both low ($R^{2}=0.05$ and $0.06$, respectively), indicating that these entropy measures capture information not captured by clustering. In contrast, LCR and VS correlated strongly with each other for both imaging and tabular datasets ($R^{2}=1.00$ and $0.90$, respectively). Fig. 7 shows results for $q=0$ and $\infty$.
Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging and (b) tabular datasets. Each point represents a dataset. All entropies are effective-number forms at $q=1$. Red line = linear regression fit; gray line = 1:1 (off the scale to the left for (a), far right). Select datasets are labeled. (c) UMAPs for labeled datasets. Gray points = unclustered elements.
[…]
When $k$ is extremely small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig. 6a), yielding a perfect correlation ($R^{2}=1.00$). When $k$ is extremely large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset (Fig. 6a), and the correlation again is perfect ($R^{2}=1.00$) (Fig. 6c). Only at intermediate $k$ values do LCR and VS provide independent information (Fig. 6b).
[…]
Figure 6: LCR vs. VS at $q=1$ at a very small half distance, the half distance with the minimum $R^{2}$, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets. (a) and (c): red line = linear regression fit; gray line = 1:1. Each point represents a dataset. (b) and (d): gray lines = $k$s for (a) and (c).
Figure 7: LCR vs. VS for $q=0$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets at their default $k$ values (11 and $\approx 2^{-1/2}$, respectively). Each point represents a dataset.
4.3 Bounds on LCR and VS
In Figs. 5 and 6, VS values were consistently higher than LCR, and both measures converge to 1 as the half‑distance decreases. It is therefore reasonable to ask whether the inequality $VS\geq LCR$ holds universally. We formalized this question as the following conjecture:
Conjecture 1.
Let $Z$ be an $n\times n$ ($n>0$) PSD similarity matrix, i.e. a PSD symmetric matrix with entries between 0 and $1$ whose diagonal entries equal $1$. Denote LCR and VS at order $q$ as $D_{q}(Z,p=\frac{1}{n})$ and $\text{VS}_{q}(Z)$, respectively, on the uniform distribution. Then for all $q\in[-\infty,+\infty]$
$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (9)
Note that in the context of this conjecture, since we are talking about the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal. We prove this for $q=2,3,$ and $\infty$ in Appendix A.2. We consider these proofs strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not $q=2.5,4$, or $27.3$. The following lemma supports this conjecture holding for all $q$:
Lemma 1.
When $Z$ is full rank but otherwise as in 1, for all $q\in[-\infty,0]$
$VS_{q}(Z)\geq n\geq D_{q}(Z,\frac{1}{n})$ (10)
Proof.
[…]
$D_{-\infty}(Z,\frac{1}{n})\geq D_{q}(Z,\frac{1}{n})$ (11)
Define $r_{i}=\sum_{j}Z_{ij}$. Then $r_{i}\geq 1$, since the entries of $Z$ are non-negative, and the diagonal entries (one of which contributes to $r_{i}$) are all $1$. Thus
$D_{-\infty}(Z,\frac{1}{n})=\frac{n}{\displaystyle\min r_{i}}\leq n$ (12)
A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$. This search used a method that successfully found a counterexample to a different inequality previously [Couch:2023pav], adding further support for Conjecture 1 being true. We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
5 Discussion
Quantification is important across most if not all complex systems. This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative. This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when every element is unique this distribution is flat. Traditional entropy is used to measure the size distribution of classes (class balance), but there is no natural way to use it to measure the dataset based on elements themselves.
5.1 Richer dataset descriptions
A popular option is to use clustering algorithms, to quantify substructure as the number of clusters (and possibly their relative sizes, which can be measured by traditional entropy). Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms also require choices. For example, HDBSCAN, which we used here, requires choosing min_cluster_size, min_samples, etc. Conceptually, the main difference is that similarity-sensitive entropy accounts for similarities both within and between clusters, whereas most clustering algorithms lack this ability: effectively, the within-cluster similarity is always 1 and the between-cluster similarity is always 0.
The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by membership strength for one or more clusters. One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster, and thereby contributing to the size of that cluster; this is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$ (a big cluster that many unique elements belong to fairly strongly is “ordinary”) (Fig. 2c). This is another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements. Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity. In this way, LCR is a richer, more accurate measure of the information in a system.
Regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, how should LCR values be interpreted? LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds the Shannon entropy. Similarity saps individual elements of some of their uniqueness.
How do LCR and VS compare with the number of clusters identified by HDBSCAN? Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Fig. 1). Indeed, the many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS. For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
In contrast, for imaging datasets, LCR and VS were typically one to two orders of magnitude lower than the number of clusters, indicating substantial similarity among clusters that hard clustering does not capture (Fig. 2c). We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure. We do not regard this as a failure of HDBSCAN; rather, we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering. However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
5.2 The effect of scaling similarity
We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑matrix values are scaled, which we illustrated using the half‑distance parameter $k$. Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward.
Why should one need to scale the similarity between elements in the first place? Should scaling be permitted at all? When similarity is derived from external knowledge or a principled model, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collates these values for all pairs. An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022]. This is the unique definition that preserves biophysical additivity; arbitrary scaling would break that relationship and thereby produce an objectively incorrect similarity matrix.
In contrast, when similarity is heuristic and not constrained by such principles, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$. For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc). Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context and/or are maximally informative about the system.
We note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency. To avoid the impression of arbitrariness, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency. It is possible that alternative values of $k$ could similarly be linked to other quantities of interest; this is left for future work.
A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$). When $k$ is small, the off‑diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element. Conversely, for large $k$ the off‑diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered. In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective‑number form of traditional entropy equals the dataset size. Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$. Other scaling schemes may exhibit the same property.
So: for heuristic similarity functions, which $k$ should be selected? Qualitatively, the answer is a value roughly midway between the extremes, but our quantitative findings indicate that the optimal choice is not obvious. We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any reasonable definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side). Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative. We also observed that even at this $k$, the $R^{2}$ between LCR and VS remained relatively high (0.680.68)—perhaps to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6.1). The dependence on $k$, however, was quite strong.
Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale). Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored. Thus to some extent, the information in a system is in the eye of the beholder.
6 Conclusion
6.1 LCR vs. VS
Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below). If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non‑trivial range (see Section 5.2) to maximize their independence and thereby their combined informativeness.
We note that, quantitatively, VS will be larger than LCR for any given $k$ and $q$. We observed this ordering empirically and proved it for a broad class of cases, outlining a proof for the remaining situations; a computational search uncovered no counter‑examples. This relationship can be rationalized by recognizing that, although VS is a similarity‑sensitive entropy, its computation includes a traditional entropy component: after incorporating similarity to derive the eigen‑elements, VS is the traditional entropy of the resulting eigenvalues; traditional entropy is always greater than or equal to similarity‑sensitive entropy.
From a practical perspective, LCR has three advantages. First, LCR is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order. Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig. 8). This performance gap may narrow or widen as algorithmic improvements are introduced.
Figure 8: Compute times in seconds on all 53 datasets for LCR (via greylock) vs. VS (via vendi-score). Imaging datasets were all 10,000 images.
Second, LCR does not require the similarity matrix to be PSD, whereas VS does. Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS. Even symmetric similarity matrices—the most common type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality. The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD. Thus, PSD‑ness is irrelevant for LCR but must be considered when applying VS.
And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components. In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class. In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$. Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research. (The greylock implementation can compute $\alpha$, $\beta$, and $\gamma$.)
The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum]. However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership. If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace). In contrast, LCR requires only the similarity matrix $Z$. Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from the element‑level similarities within the classes. In principle, the VS formulation permits two classes that contain exactly the same individuals to nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains where traditional entropy is a mainstay. We anticipate that the deeper understanding of their relationship and properties as presented here will aid such investigations.
Appendix A Appendix
A.1 Python code for finding counterexamples to $VS\geq LCR$
import numpy as np
import pandas as pd
from scipy.optimize import minimize
from tqdm import tqdm, trange
from multiprocessing import Pool
n_workers = 96
def uniform_div(Z,q):
    p = (1.0/Z.shape[0])*np.ones_like(Z[0])
    Zp = Z@p
    if q==1:
        h1 = -np.sum(p*np.log(Zp))
        return np.exp(h1)
    else:
        X = np.sum(p*(Zp**(q-1)))
        h = (1/(1-q)) * np.log(X)
        return np.exp(h)
def vendi_score(Z,q):
    evals = np.linalg.eigvalsh(Z/Z.shape[0])
    evals = np.array([e for e in evals if e>0]) # keep only the support
    if q==1:
        h1 = -np.sum(evals*np.log(evals))
        return np.exp(h1)
    else:
        X = np.sum(evals**q)
        h = (1/(1-q)) * np.log(X)
        return np.exp(h)
def func(x, n, r, q):
    Z = x_to_Z(x, n, r)
    return vendi_score(Z, q) - uniform_div(Z, q)
def x_to_Z(x, n, r):
    # map from reals to positive reals, reshape
    w = (x**2).reshape(r, n)
    # normalize rows
    w = w/np.sqrt((w**2).sum(0))
    Z = w.transpose()@w
    return Z
def find_min(n, r, q, scale):
    args = (n, r, q)
    x0 = scale*np.random.normal(0, 1, n*r)
    min_obj = minimize(func, x0, args)
    ret = dict(min_obj)
    ret.update({’x0’: x0, ’Z0’: x_to_Z(x0, n, r), ’Z’: x_to_Z(min_obj.x, n, r)})
    return ret
qlist = [-10, -2.7, -2, -1, -0.3, 0, 0.3, 1, 2.7, 3.4, 4, 5, 2*np.pi, 7, 8, 9.1, 10]
sizes = [3, 5, 10, 20]
scales = [0.1, 1, 2, 10]
replicates = [i for i in range(1,6)]
combos = [{’q’: q, ’n’: n, ’r’: r, ’rep’: rep, ’scale’: s} \
for q in qlist for n in sizes for s in scales \
for rep in replicates \
for r in set([x for x in [2, n//2, n] if x>1])]
def map_func(combo):
    q = combo[’q’]
    r = combo[’r’]
    n = combo[’n’]
    rep = combo[’rep’]
    scale = combo[’scale’]
    row = combo
    try:
        row.update(find_min(n, r, q, scale))
        return row
    except np.linalg.LinAlgError:
        return row
results = []
np.random.shuffle(combos)
n_batches = len(combos)//n_workers
if n_batches*n_workers< len(combos):
    n_batches +=1
pool = Pool(n_workers)
for batch in trange(n_batches):
    ix_start = batch*n_workers
    ix_end = ix_start + n_workers
    rows = pool.map(map_func, combos[ix_start:ix_end])
    results = results + rows
    results_df = pd.DataFrame(results)
    results_df.to_pickle(’../data/vendi_inequality_counterexample_search.pkl’)
A.2 Proof that $VS\geq LCR$ for $q\in\{2,3,\infty\}$
Theorem 1.
Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then when $q\in\{2,3,\infty\}$,
$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (13)
Proof.
We will consider the cases separately. For convenience, in what follows, denote by $\lambda_{i}$ the eigenvalues of $Z$ and $r_{i}$ the row sums of $Z$, i.e. $r_{i}=\sum_{j}Z_{ij}$.
At $q=2$,
[…]
However, since $0\leq Z_{ij}\leq 1\forall(i,j)$, it follows that $Z_{i,j}^{2}\leq Z_{ij}$ and so $\sum_{i,j}Z_{i,j}^{2}\leq\displaystyle\sum_{i,j}Z_{ij}$. Therefore,
$\frac{n^{2}}{\sum_{i,j}Z^{2}_{i,j}}\geq\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ (15)
However, $\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ is simply $D_{2}(Z,\frac{1}{n})$, proving the $q=2$ case.
For $q=3$, we have that
[…]
$=\sum_{i,j,k}Z_{ik}Z_{kj}=\sum_{k}\left(\sum_{i}Z_{ki}\right)\left(\sum_{j}Z_{kj}\right)=\sum_{i}r_{i}^{2}.$ (17)
This gives
[…]
proving the $q=3$ case.
For $q=\infty$, we invoke the Gershgorin circle inequality. We first note that
$\mathrm{VS}_{\infty}(Z)$ $=$ $\frac{n}{\lambda_{\text{max}}}$ (19)
$D_{\infty}(Z,\frac{1}{n})$ $=$ $\frac{n}{r_{\text{max}}}$ (20)
where $\lambda_{\text{max}}$ is the largest eigenvalue of $Z$, i.e. $\lambda_{\text{max}}=\displaystyle\max_{i}\lambda_{i}$, and likewise $r_{\text{max}}=\displaystyle\max_{i}r_{i}$. However, by the Gershgorin circle inequality, every eigenvalue of $Z$ must fall within a Gershgorin interval, i.e. in one of the intervals $Z_{ii}-\sum_{j\neq i}Z_{i,j},Z_{ii}+\sum_{j\neq i}Z_{i,j}]$. As such, noting that $Z_{ii}+\sum_{j\neq i}Z_{i,j}=\sum_{j}Z_{i,j}=r_{i}$, we must have that every eigenvalue $\lambda_{\mu}$ of $Z$ is bounded by $\lambda_{\mu}\leq r_{\text{max}}$. But this means that
$\mathrm{VS}_{\infty}(Z)=\frac{n}{\lambda_{\text{max}}}\geq\frac{n}{r_{\text{max}}}=D_{\infty}(Z,\frac{1}{n})$ (21)
A.3 Approach to a proof of $VS\geq LCR$ for all $q$
A proof for 1 at all $q$ except for $1$ (which could perhaps be filled in using continuity considerations) would follow from the following:
Conjecture 2.
Let $Z$ be as in 1, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for $q>1$,
[…]
Likewise, for $q<1$,
[…]
At $q>1$, 2 would in turn be proven by the following conjecture:
Conjecture 3.
Let $M$ be any $n\times n$ symmetric matrix whose entries obey $0\leq M_{ij}\leq 1$, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for any $\alpha\geq 0$
$\mathbf{1}^{\intercal}M^{\alpha}\mathbf{1}\leq\mathbf{1}^{\intercal}\left(M\mathbf{1}\right)^{\alpha}$ (24)
A.4 Other bounds
Beyond the conjectures upper bound of uniform diversity-with-similarity by Vendi score, we are able to prove the following bound in the opposite direction:
Theorem 2.
Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then
$\mathrm{VS}_{1}(Z)\leq\frac{1}{n}\left(\sum_{i,j}Z_{ij}\right)D_{1}(Z,\frac{1}{n})$ (25)
Proof.
We begin by constructing the diagonal matrix $W$ according to
$W_{ii}$ $=$ $\frac{1}{T}\sum_{j}Z_{ij}$ (26)
$T$ $=$ $\sum_{i,j}Z_{ij}$ (27)
[…]
tr[$\tr\left[\frac{Z}{n}\left(\log\frac{Z}{n}-\log W\right)\right]\geq 0$ (28)
$\Rightarrow$
[…]
However, since $W$ is diagonal we have that
[…]
$=$ $-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{T}$ (31)
$=$ $-\sum_{i}\frac{1}{n}\left(\log\frac{\sum_{j}Z_{ij}}{n}-\log\frac{T}{n}\right)$ (32)
$=$ $\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (33)
As such, we have that
$\log\mathrm{VS}_{1}(Z)\leq\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (34)
Exponentiating both sides, and recognizing that $e^{-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}}=D_{1}(Z,\frac{1}{n})$, we have that
$\mathrm{VS}_{1}(Z)\leq\frac{T}{n}D_{1}(Z,\frac{1}{n})$ (35)
and so the theorem is proved.

### v2

@@
-Which Similarity-Sensitive Entropy?
+Which Similarity-Sensitive Entropy (S-entropy)?
 Abstract
@@
 Shannon entropy and other traditional entropy measures capture only the information encoded in the frequencies of a system’s elements.
-Recently, Leinster, Cobbold, and Reeve (LCR) introduced a method that also captures the rich information encoded in the similarities and differences among elements, yielding similarity-sensitive entropy.
-More recently, the Vendi score (VS) was introduced as an alternative, raising the question of how LCR and VS compare, and which is preferable.
+Recently, Leinster, Cobbold, and Reeve (LCR) introduced a method that also captures the far richer additional information encoded in the similarities and differences among elements, yielding similarity-sensitive entropy (S-entropy).
+More recently, the Vendi score (VS) was introduced as an alternative, raising the practical questions of how LCR and VS compare and which one to use.
 Here we address these questions conceptually, analytically, and experimentally, using 53 machine-learning datasets.
-We show that LCR and VS can differ by orders of magnitude and can capture complementary information about a system, except in limiting cases.
+We show that LCR and VS values can differ by orders of magnitude and can capture complementary information about a system, except in limiting cases.
 We demonstrate that both LCR and VS depend on how similarities are scaled and introduce the concept of “half distance” to parameterize this dependence.
-We prove that VS provides an upper bound on LCR for several values of the Rényi-Hill order parameter and conjecture that this bound holds for all values.
+We prove that VS provides an upper bound on LCR for several values of the Rényi-Hill order parameter and present evidence that this bound holds for all values.
 We conclude that VS is preferable only when interpreting elements as linear combinations of a more fundamental set of “ur-elements” or when the system or dataset possesses a quantum-mechanical character.
-In the broader circumstance where one seeks simply to capture the rich information encoded by similarity, LCR is favored; nevertheless, for certain half-distances the two methods can complement each other.
+In the broader circumstance where one seeks simply to capture the rich information encoded by similarity as part of one’s entropy calculation, LCR is favored; nevertheless, for certain half-distances the two methods can complement each other.
 1 Introduction
@@
 1).
-Figure 1: The concepts of element, frequency distribution, traditional entropy (at $q=1$), similarity function, similarity measure ($z_{i}j$, with examples), similarity matrix ($Z$), and similarity-sensitive entropy (here, LCR, also at $q=1$).
+Figure 1: The concepts of element, frequency distribution, traditional entropy (at $q=1$), similarity function, similarity measure ($z_{i}j$, with examples), similarity matrix ($Z$), and S-entropy (here, LCR, also at $q=1$).
 Entropy values are expressed in effective-number form, i.e.
 in units of effective number of images present in the dataset.
-Different similarity measures can be chosen (see Section 5.2); in this illustrative example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
-Incorporating similarities into the calculation of entropy results in so-called similarity-sensitive entropy.
+Different similarity measures can be chosen (see Section 5.2); in this example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
+Incorporating similarities into the calculation of entropy results in so-called similarity-sensitive entropy or S-entropy.
 This was first accomplished by Leinster and Cobbold [leinster2012measuring] and extended by Reeve and colleagues [reeve2014partition], in a framework we refer to here as LCR [leinsterEntropyDiversityAxiomatic2020].
@@
 similarity-insensitive) entropies because their frequency distributions are both flat.
-Similarity-sensitive entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [arnaoutFutureBloodTesting2021, aroraRepertoirescaleMeasuresAntigen2022], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat and therefore traditional entropies less informative.
-More recently, an alternative form of similarity-sensitive entropy has been described, consisting of the Vendi score [friedman2022vendi, pasarkar2025vendiscope] and its variants [pasarkar2023cousins]—hereafter collectively VS—separate from LCR.
-This raises the practical question of which form of similarity-sensitive entropy to choose for a given situation: LCR or VS.
+S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [arnaoutFutureBloodTesting2021, aroraRepertoirescaleMeasuresAntigen2022], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat and therefore traditional entropies less informative.
+More recently, an alternative form of S-entropy has been described, consisting of the Vendi score [friedman2022vendi, pasarkar2025vendiscope] and its variants [pasarkar2023cousins]—hereafter collectively VS—separate from LCR.
+This raises the practical question of which form of S-entropy to choose for a given situation: LCR or VS.
 The current work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets used for ML, and comparing their values.
@@
 Entropies fall into two classes: traditional/similarity‑insensitive and similarity‑sensitive.
-Although one often sees reference to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), entropies are actually families of measures; individual members are distinguished by how they weight frequencies and, for similarity-sensitive entropy, by their similarity matrix (see Table 1).
-2.1 Traditional entropy (similarity-insensitive)
+Although one often sees reference to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), entropies are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix (see Table 1).
+2.1 Traditional, similarity-insensitive entropy
 […]
 Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the system [jost2006entropy].
-Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types, while larger $q$ give greater weight to common types.
+Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements.
 $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form.
-Hill demonstrated that several other familiar statistics are special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [hill1973diversity].
+Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [hill1973diversity].
 For this reason, Hill’s formulation is often described as a unifying framework.
@@
 where $\mathbf{p}$ denotes the frequency distribution of unique elements.
-Other generalizations, such as Tsallis entropy [tsallis1988possible], also exist.
-What these traditional entropies all have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
-2.2 Similarity-sensitive entropy
+Other generalizations such as Tsallis entropy [tsallis1988possible] also exist.
+What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
+2.2 Similarity-sensitive entropy (S-entropy)
 2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
-The LCR framework [leinsterEntropyDiversityAxiomatic2020] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropy does not.
+LCR [leinsterEntropyDiversityAxiomatic2020] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropy does not.
 A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$.
 The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1.
-The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive Rényi entropies $H_{\alpha}^{Z}$ and are given by
+The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
 $D_{q}^{Z}(\mathbf{p};Z)=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}$ (2)
@@
   • attribute diversity [chaoAttributediversityApproachFunctional2019] when interpreting the system as a set of attribute contributions instead of frequencies; and
-  • similarity-sensitive or similarity-aware diversity [nguyen2023greylock] more generally.
-LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become uninformative.
-Representative applications in the life sciences include high‑throughput immunology (immunomes) [aroraRepertoirescaleMeasuresAntigen2022], microbiome research (metagenomics) [nguyen2023greylock], and medical imaging [couch2024beyond].
-specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g.
+  • similarity-sensitive or similarity-aware diversity more generally [nguyen2023greylock].
+LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become uninformative (Fig.
+1).
+Representative applications in the life sciences include high‑throughput immunology (immunomes) [aroraRepertoirescaleMeasuresAntigen2022], microbiome research (metagenomics) [nguyen2023greylock], and medical imaging [couch2024beyond] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g.
 images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [couch2024beyond].
@@
 VS constitutes a related but separate class of similarity‑sensitive entropy measures.
-Like LCR, VS entropies are functions of a similarity matrix, but the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element.
-We can define $Z_{n}$ in terms of $Z$ as
+Like LCR, VS entropies are functions of a similarity matrix, but now the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element.
+(When all elements are unique, $Z$ and $Z_{n}$ are the same.) We can define $Z_{n}$ in terms of $Z$ as
 $\left(Z_{n}\right)_{i,j}=Z_{s(i),s(j)}.$ (3)
-where $s(i)$ is the unique element of which the $i^{\mathbf{th}}$ overall element is an example.
+where $s(i)$ is the unique element of which the $i^{\text{th}}$ overall element is an instance or example.
 The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\lambda_{i}$ of $Z_{n}/n$.
@@
 $\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\right)$ (4)
-VS is to the similarity matrix what the von Neumann entropy is to the quantum density matrix.
+VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix.
 Replacing the Shannon term in Eq.
@@
 3.1 Datasets and definitions
-The following imaging and tabular datasets were downloaded for benchmarking LCR and VS:
+The following imaging and tabular datasets were used for benchmarking LCR and VS:
   • Image datasets.
-22 large ($\geq 10,000$-image) standard vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [yangMedMNISTClassificationDecathlon2021a]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [visipediaInat_comp2017READMEmd]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [mei2022radimagenet], the NIH ChestXRay dataset [wangChestXRay8HospitalScaleChest2017]; COCO (Common Objects in Context) [linMicrosoftCOCOCommon2015]; and MIDRC COVIDx CXR-4 [wuCOVIDxCXR4Expanded2023].
+22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [yangMedMNISTClassificationDecathlon2021a]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [visipediaInat_comp2017READMEmd]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [mei2022radimagenet], the NIH ChestXRay dataset [wangChestXRay8HospitalScaleChest2017]; COCO (Common Objects in Context) [linMicrosoftCOCOCommon2015]; and MIDRC COVIDx CXR-4 [wuCOVIDxCXR4Expanded2023].
 For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
   • Tabular datasets.
-After selecting for a maximum of 30 columns and 200,000 rows, the 31 most popular datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [lucieImprovedPython2024].
+After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [lucieImprovedPython2024].
 For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
-Definitions.
-The terms “system” and “dataset” are used interchangeably.
+Definitions (Fig.
+1).The terms “system” and “dataset” are used interchangeably.
 The elements of image and tabular datasets are single images and single rows, respectively.
@@
 7 equivalent to the root-mean-square distance (RMSD).
+The default $k$ for tabular datasets was $1$.
 3.3 Entropy calculations
-The greylock Python package [nguyen2023greylock] was used to compute LCR as the $\gamma$ diversity; this is the effective number form of the corresponding entropy (see Section 6).
-The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with VS.
+The greylock Python package [nguyen2023greylock] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1).
+The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2).
 VS was calculated using the vendi‑score package [vendi-score].
 This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR.
-All computations were performed on a workstation equipped with multi‑core CPUs (greylock leverages parallelization).
+Similarity matrices were calculated on a GPU-enabled workstation equipped with multi‑core CPUs.
+LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
 3.4 Clustering
@@
 A negative value indicates a counterexample.
-To meet the assumptions of this conjecture, we constructed $Z$ as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$.
+To meet the assumptions of this conjecture, $Z$ was constructed as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$.
 To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative.
@@
 We minimized the difference $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for each combination of $q\in\{-10$, 2.7-2.7, 2-2, 1-1, 0.3-0.3, 0, $0.3$, 11, $2.7$, $3.4$, 44, 55, $2\pi$, $7$, 88, $9.1$, 10}10\}, $n\in\{3,5,10,20\}$, $r\in\{2,\lfloor n/2\rfloor,n\}$ if $n>3$ and $r\in\{2,3\}$ for $n=3$ (to avoid $r=1$).
-We also vary the distribution from which the initial guess fed to scipy.optimize was drawn.
+We also varied the distribution from which the initial guess fed to scipy.optimize was drawn.
 In total, 4420 minimization were done.
 The full Python code for performing these minimizations can be found in section A.1
+3.6 Speed tests
+Run times were calculated using the Python timeit package with 5-fold replications and measured on a 2024 M4 Apple Mac mini with 64GB unified memory.
 4 Results
@@
 Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
+Figure 2: The first 10,000 images of the MNIST digits dataset colored (a) by digit (labeled using representative images), (b) by HDBSCAN cluster, and (c) by similarity to one of the images (a “9”).
+The effective number of images in this dataset is 12.5 by LCR and 95.9 by VS.
 Fig.
@@
 with zero pairwise similarity).
-This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the similarity-sensitive entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease the entropy relative to a scenario in which all 10 digit classes were entirely distinct.
+This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease the entropy relative to a scenario in which all 10 digit classes were entirely distinct.
 In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements.
-However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑images (or “ur‑images”), each actual image being a linear combination of these ur-images.
-Figure 3 displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy.
+However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑ or “ur”‑images, each actual image in the dataset being a linear combination of these.
+Fig.
+3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig.
+3b).
 Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits.
@@
 The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small.
-Figure 2: The first 10,000 images of the MNIST digits dataset colored (a) by digit (labeled using representative images), (b) by HDBSCAN cluster, and (c) by similarity to one of the images (a “9”).
-The effective number of images in this dataset is 12.5 by LCR and 95.9 by VS.
 Figure 3: (a) Top 100 eigenimages for the MNIST digits dataset and (b) their eigenvalues (in the gray region), alongside the full eigenvalue spectrum. Note the log axes.
-Figure 4: LCR and VS for (a) imaging and (b) tabular datasets, sorted by VS ($q=1$; default $k$s).
 Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets.
+Figure 4: LCR and VS for (a) imaging and (b) tabular datasets, sorted by VS ($q=1$; default values of $k$ for each dataset type).
 4.2 Trends with LCR and VS
 Fig.
-5 presents comparisons among LCR, VS, and the number of HDBSCAN clusters across all datasets, with imaging and tabular datasets evaluated separately.
+5 presents comparisons among LCR, VS, and the number of HDBSCAN clusters across all datasets (for default $k$ at $q=1$; Figs.
+6 and 7 investigate other $k$s and $q$s), with imaging and tabular datasets evaluated separately.
 As observed for MNIST, both LCR and VS were generally larger than the cluster count.
@@
 Select datasets are labeled.
-(c) UMAPs for labeled datasets.
+(c) UMAPs for labeled imaging datasets.
 Gray points = unclustered elements.
 […]
-When $k$ is extremely small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig.
-6a), yielding a perfect correlation ($R^{2}=1.00$).
-When $k$ is extremely large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset (Fig.
-6a), and the correlation again is perfect ($R^{2}=1.00$) (Fig.
-6c).
+When $k$ is very small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig.
+6a and c, left), yielding a perfect correlation ($R^{2}=1.00$).
+When $k$ is very large, all off-diagonal $z_{ij}$ approach 0, making every element essentially unique; consequently, both LCR and VS converge to the number of elements in the dataset, and the correlation again is perfect ($R^{2}=1.00$) (Fig.
+6a and c, right).
 Only at intermediate $k$ values do LCR and VS provide independent information (Fig.
-6b).
+6a and c, middle).
 […]
 Figure 6: LCR vs.
-VS at $q=1$ at a very small half distance, the half distance with the minimum $R^{2}$, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets.
+VS at $q=1$ at a very small half distance, an intermediate half-distance for which the correlation is low, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets.
+For tabular datasets, the intermediate half-distance is the one simultaneously minimizing the $R^{2}$, the correlation, and maximizing the covariance at $q=1$.
+For image datasets, the intermediate half-distance does not coincide with the global minimum of the correlation at $q=1$, because VSs and LCRs already come very close to 10,000 at that global minimum, so we chose a smaller value.
 (a) and (c): red line = linear regression fit; gray line = 1:1.
 Each point represents a dataset.
-(b) and (d): gray lines = $k$s for (a) and (c).
+(b) and (d): vertical gray lines = $k$s for (a) and (c).
 Figure 7: LCR vs. VS for $q=0$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets at their default $k$ values (11 and $\approx 2^{-1/2}$, respectively). Each point represents a dataset.
@@
 5 Discussion
-Quantification is important across most if not all complex systems.
+Quantifying the information contained in a system is essential for virtually any complex system.
 This includes ML datasets, where the fact that most elements are generally unique—no repeated images (in image datasets) or rows (in tabular datasets)—renders traditional entropy (Shannon entropy and the other Rényi entropies) uninformative.
-This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when every element is unique this distribution is flat.
-Traditional entropy is used to measure the size distribution of classes (class balance), but there is no natural way to use it to measure the dataset based on elements themselves.
-5.1 Richer dataset descriptions
-A popular option is to use clustering algorithms, to quantify substructure as the number of clusters (and possibly their relative sizes, which can be measured by traditional entropy).
-Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms also require choices.
-For example, HDBSCAN, which we used here, requires choosing min_cluster_size, min_samples, etc.
-Conceptually, the main difference is that similarity-sensitive entropy accounts for similarities both within and between clusters, whereas most clustering algorithms lack this ability: effectively, the within-cluster similarity is always 1 and the between-cluster similarity is always 0.
-The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by membership strength for one or more clusters.
-One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster, and thereby contributing to the size of that cluster; this is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$ (a big cluster that many unique elements belong to fairly strongly is “ordinary”) (Fig.
+This is because traditional entropy measures information encoded by the shape of the elements’ frequency distribution, but when the elements each appear only once, this distribution is flat (Fig.
+1): there is no additional information there.
+Traditional entropy is used to measure the distribution of class sizes, which often vary; this is class balance.
+However, there is no obvious way to use class balance to measure the the information contained within or between classes, which is a property of the elements themselves.
+5.1 Richer measures of the information that ML datasets contain
+One way to get at this missing information is to use clustering algorithms, to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy).
+Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices.
+For example, HDBSCAN requires choosing min_cluster_size, min_samples, etc.
+Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
+The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters.
+One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster and thereby contributing to the size of that cluster.
+This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster that many unique elements belong to fairly strongly is “ordinary” (Fig.
 2c).
-This is another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
+This perspective provides another way of seeing that LCR is the entropy of a system after accounting/adjusting/correcting for the similarity among its elements.
 Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity.
 In this way, LCR is a richer, more accurate measure of the information in a system.
-Regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, how should LCR values be interpreted?
-LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds the Shannon entropy.
-Similarity saps individual elements of some of their uniqueness.
+Note that, regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds Shannon entropy.
+Similarity saps individual elements of some of their uniqueness, but in doing so, it reveals differences between datasets that traditional entropy cannot (Fig.
+1).
 How do LCR and VS compare with the number of clusters identified by HDBSCAN?
-Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Fig.
-1).
-Indeed, the many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS.
+Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs.
+1, 5c).
+Indeed, many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS.
 For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
@@
 We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure.
-We do not regard this as a failure of HDBSCAN; rather, we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering.
+We do not regard this as a failure of HDBSCAN; we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering.
 However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
-5.2 The effect of scaling similarity
-We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑matrix values are scaled, which we illustrated using the half‑distance parameter $k$.
-Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward.
+5.2 The effect of scaling the similarity measure/similarity matrix
+We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$.
+Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward in the middle, while maintaining the anchoring at each end (at 0 and 1).
 Why should one need to scale the similarity between elements in the first place?
 Should scaling be permitted at all?
-When similarity is derived from external knowledge or a principled model, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collates these values for all pairs.
+When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs.
 An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022].
-This is the unique definition that preserves biophysical additivity; arbitrary scaling would break that relationship and thereby produce an objectively incorrect similarity matrix.
+This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
 In contrast, when similarity is heuristic and not constrained by such principles, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$.
@@
 Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context and/or are maximally informative about the system.
-We note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency.
-To avoid the impression of arbitrariness, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency.
-It is possible that alternative values of $k$ could similarly be linked to other quantities of interest; this is left for future work.
+Such choices arise often in measurement.
+For example, we note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency.
+Lest choices feel too arbitrary, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency.
+It is possible that alternative values of $k$ could similarly be linked to other quantities of interest.
+Investigating this possibility is left for future work.
 A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$).
@@
 Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$.
-Other scaling schemes may exhibit the same property.
-So: for heuristic similarity functions, which $k$ should be selected?
-Qualitatively, the answer is a value roughly midway between the extremes, but our quantitative findings indicate that the optimal choice is not obvious.
-We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any reasonable definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side).
+Other scaling approaches may exhibit the same property.
+So: for heuristic similarity functions, which $k$ should one choose?
+Qualitatively, the answer is a value roughly midway between the extremes (Fig.
+6).
+However, our quantitative findings indicate that the optimal choice is not obvious.
+We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side).
 Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative.
-We also observed that even at this $k$, the $R^{2}$ between LCR and VS remained relatively high (0.680.68)—perhaps to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6.1).
+We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$.
+Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6).
 The dependence on $k$, however, was quite strong.
-Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale).
+Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale).
 Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored.
@@
 6 Conclusion
-6.1 LCR vs. VS
+So: LCR or VS?
 Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below).
@@
 From a practical perspective, LCR has three advantages.
-First, LCR is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order.
+First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order.
 Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig.
 8).
-This performance gap may narrow or widen as algorithmic improvements are introduced.
-Figure 8: Compute times in seconds on all 53 datasets for LCR (via greylock) vs. VS (via vendi-score). Imaging datasets were all 10,000 images.
+This performance gap may narrow or widen as codebases evolve.
+Figure 8: Compute times for all 53 datasets for LCR (using the greylock Python package [nguyen2023greylock]) vs.
+VS (using vendi-score).
+Imaging datasets were all 10,000 images.
+Smaller datasets were faster.
+Note the different timescales.
 Second, LCR does not require the similarity matrix to be PSD, whereas VS does.
 Similarity matrices need not be PSD; a perfectly reasonable set of similarity measurements can produce a matrix that fails the PSD condition, which would complicate the use of VS.
-Even symmetric similarity matrices—the most common type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality.
-The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD.
-Thus, PSD‑ness is irrelevant for LCR but must be considered when applying VS.
-And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components.
+Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality.
+The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [leinster2013magnitude]).
+Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
+And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] (Table 1) is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components.
 In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class.
@@
 Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research.
-(The greylock implementation can compute $\alpha$, $\beta$, and $\gamma$.)
+(The greylock implementation computes $\alpha$, $\beta$, and $\gamma$.)
 The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum].
 However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership.
-If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace).
+If there are $n$ total elements and $k$ classes, the VS decomposition demands either an $nk\times nk$ matrix or an $n\times k\times n\times k$ 4-dimensional tensor (which can be reduced to $n\times n$ and $k\times k$ matrices by taking a partial trace).
 In contrast, LCR requires only the similarity matrix $Z$.
-Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from the element‑level similarities within the classes.
-In principle, the VS formulation permits two classes that contain exactly the same individuals to nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
-In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains where traditional entropy is a mainstay.
-We anticipate that the deeper understanding of their relationship and properties as presented here will aid such investigations.
+Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes.
+In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
+In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay.
+We anticipate that the deeper understanding of their relationship and properties through work such as that presented here will aid such investigations.
+7 Acknowledgements
+This work was supported by the Gordon and Betty Moore Foundation and by the NIH under grants R01HL150394, R01HL150394-SI, R01AI148747, and R01AI148747-SI.
 Appendix A Appendix

### v3

@@
 Abstract
-A canonical step in quantifying a system is to measure its entropy.
-Shannon entropy and other traditional entropy measures capture only the information encoded in the frequencies of a system’s elements.
-Recently, Leinster, Cobbold, and Reeve (LCR) introduced a method that also captures the far richer additional information encoded in the similarities and differences among elements, yielding similarity-sensitive entropy (S-entropy).
-More recently, the Vendi score (VS) was introduced as an alternative, raising the practical questions of how LCR and VS compare and which one to use.
-Here we address these questions conceptually, analytically, and experimentally, using 53 machine-learning datasets.
-We show that LCR and VS values can differ by orders of magnitude and can capture complementary information about a system, except in limiting cases.
-We demonstrate that both LCR and VS depend on how similarities are scaled and introduce the concept of “half distance” to parameterize this dependence.
-We prove that VS provides an upper bound on LCR for several values of the Rényi-Hill order parameter and present evidence that this bound holds for all values.
-We conclude that VS is preferable only when interpreting elements as linear combinations of a more fundamental set of “ur-elements” or when the system or dataset possesses a quantum-mechanical character.
-In the broader circumstance where one seeks simply to capture the rich information encoded by similarity as part of one’s entropy calculation, LCR is favored; nevertheless, for certain half-distances the two methods can complement each other.
+Shannon entropy is not the only entropy that is relevant to machine‑learning datasets, nor possibly even the most important one.
+Traditional entropies such as Shannon entropy capture information represented by elements’ frequencies but not the richer information encoded by their similarities and differences.
+Capturing the latter requires similarity‑sensitive entropy (S‑entropy).
+S‑entropy can be measured using either the recently developed Leinster–Cobbold–Reeve framework (LCR) or the newer Vendi score (VS).
+This raises the practical question of which one to use: LCR or VS.
+Here we address this question conceptually, analytically, and experimentally, using 53 large and well-known imaging and tabular datasets.
+We find that LCR and VS values can differ by orders of magnitude and are complementary, except in limiting cases.
+We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half‑distance” to parameterize this dependence.
+We prove that VS provides an upper bound on LCR for several values of the Rényi–Hill order parameter and present evidence that this bound holds for all values.
+We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur‑elements” or when the system that the dataset describes has a quantum‑mechanical character.
+In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, LCR is favored; nevertheless, for certain half‑distances the two methods can complement each other.
 1 Introduction
-Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across science, engineering, commerce, and law [leinster2016maximizing, Shannon1951, Montemurro1994, tawilCompetitionCaliforniasMediCal2022].
-Traditional entropy, as formulated by Shannon [shannon1948mathematical] and generalized by Rényi [renyiMeasuresEntropyInformation1961] (and later Tsallis [tsallis1988possible]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits or nats required to encode the shape of the distribution, with smaller values for more skew distributions and larger values for flatter ones (Fig.
+Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across not only science and engineering but in commerce and law [15, 29, 22, 31].
+Traditional entropy, as formulated by Shannon [30] and generalized by Rényi [28] (and later Tsallis [32]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig.
 1).
-These measures rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig.
+These entropies rely solely on frequencies and therefore ignore a rich additional source of information about the system: the pairwise similarities and differences among its elements (Fig.
 1).
@@
 Different similarity measures can be chosen (see Section 5.2); in this example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
-Incorporating similarities into the calculation of entropy results in so-called similarity-sensitive entropy or S-entropy.
-This was first accomplished by Leinster and Cobbold [leinster2012measuring] and extended by Reeve and colleagues [reeve2014partition], in a framework we refer to here as LCR [leinsterEntropyDiversityAxiomatic2020].
+Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or S-entropy.
+This was first accomplished by Leinster and Cobbold [14] and extended by Reeve and colleagues [27], in a framework we refer to as LCR [17].
 By incorporating similarity, LCR can differentiate between systems that have identical frequency distributions but whose elements vary in how alike they are (Fig.
 1).
-This capability is valuable in many domains [nguyen2023greylock], especially in machine learning (ML) [couch2024beyond, couch2025x] (for example, to achieve state-of-the-art performance more efficiently [chinnENRICHingMedicalImaging2023]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e.
-similarity-insensitive) entropies because their frequency distributions are both flat.
-S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [arnaoutFutureBloodTesting2021, aroraRepertoirescaleMeasuresAntigen2022], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat and therefore traditional entropies less informative.
-More recently, an alternative form of S-entropy has been described, consisting of the Vendi score [friedman2022vendi, pasarkar2025vendiscope] and its variants [pasarkar2023cousins]—hereafter collectively VS—separate from LCR.
-This raises the practical question of which form of S-entropy to choose for a given situation: LCR or VS.
-The current work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets used for ML, and comparing their values.
+This capability is valuable in many domains [23], especially in machine learning (ML) [6, 7] (for example, to achieve state-of-the-art performance more efficiently [5]), where datasets generally consist of all-unique elements, meaning any two datasets that are the same size will have the same traditional (i.e.
+similarity-insensitive) entropies because their frequency distributions are both flat, even though they may differ dramatically in composition.
+S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [1, 2], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
+More recently, an alternative form of S-entropy to LCR has been described, consisting of the Vendi score [9, 26] and its variants or “cousins” [25]—hereafter collectively VS.
+This alternative raises the practical question of which form of S-entropy to choose for a given system or dataset: LCR or VS.
+The present work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets commonly used for ML research, and comparing their values.
 2 Related Work
-Entropies fall into two classes: traditional/similarity‑insensitive and similarity‑sensitive.
-Although one often sees reference to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), entropies are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix (see Table 1).
+Entropies fall into two classes: (i) traditional, i.e.
+similarity‑insensitive, and (ii) similarity‑sensitive, a.k.a.
+S-entropy.
+Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix as well (see Table 1).
 2.1 Traditional, similarity-insensitive entropy
 […]
-Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the system [jost2006entropy].
+Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the dataset or system [13].
 Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements.
 $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form.
-Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [hill1973diversity].
+Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [12].
 For this reason, Hill’s formulation is often described as a unifying framework.
@@
 $D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1.\end{cases}$ (1)
-where $\mathbf{p}$ denotes the frequency distribution of unique elements.
-Other generalizations such as Tsallis entropy [tsallis1988possible] also exist.
+$\mathbf{p}$ denotes the frequency distribution of unique elements.
+Other generalizations such as Tsallis entropy [32] also exist.
 What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
-2.2 Similarity-sensitive entropy (S-entropy)
+2.2 S-entropy
 2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
-LCR [leinsterEntropyDiversityAxiomatic2020] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropy does not.
+LCR [17] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropies do not.
 A similarity matrix $Z$ is introduced, with entries $z_{ij}\in[0,1]$ that quantify the similarity between elements $i$ and $j$.
@@
 When this average is large, element $i$ is considered ordinary, making Eq.
-2 interpretable as the average “ordinariness” [leinster2012measuring] across all elements.
+2 interpretable as the average “ordinariness” [14] across all elements.
 $D_{q}^{Z}{(\mathbf{p})}$ appears widely in the recent literature, where it is variously known as:
-  • phylogenetic diversity [chao2010phylogenetic] in the special case where $z_{ij}$ forms an ultrametric, typically derived from a phylogenetic tree;
-  • functional diversity [chaoAttributediversityApproachFunctional2019] when the similarity pertains to elements’ function, for example the binding similarity between pairs of antibodies or TCRs [aroraRepertoirescaleMeasuresAntigen2022];
-  • attribute diversity [chaoAttributediversityApproachFunctional2019] when interpreting the system as a set of attribute contributions instead of frequencies; and
-  • similarity-sensitive or similarity-aware diversity more generally [nguyen2023greylock].
-LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become uninformative (Fig.
+  • phylogenetic diversity [3] in the special case where $z_{ij}$ forms an ultrametric, typically derived from a phylogenetic tree;
+  • functional diversity [4] when the similarity pertains to elements’ function, for example the binding similarity between pairs of antibodies or TCRs [2];
+  • attribute diversity [4] when interpreting the system as a set of attribute contributions instead of frequencies; and
+  • similarity-sensitive or similarity-aware diversity more generally [23], for which “S-entropy” is a convenient shorthand.
+LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig.
 1).
-Representative applications in the life sciences include high‑throughput immunology (immunomes) [aroraRepertoirescaleMeasuresAntigen2022], microbiome research (metagenomics) [nguyen2023greylock], and medical imaging [couch2024beyond] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g.
-images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [couch2024beyond].
-2.2.2 The Vendi score and its “cousins” (VS)
+Representative applications in the life sciences include high‑throughput immunology (immunomes) [2], microbiome research (metagenomics) [23], and medical imaging [6] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g.
+images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [6].
+2.2.2 The Vendi score and its variants (VS)
 VS constitutes a related but separate class of similarity‑sensitive entropy measures.
@@
 Replacing the Shannon term in Eq.
-4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [pasarkar2023cousins] of the Vendi score:
+4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [25] of the Vendi score:
 $\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\lambda_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
@@
   • Image datasets.
-22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [yangMedMNISTClassificationDecathlon2021a]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [visipediaInat_comp2017READMEmd]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [mei2022radimagenet], the NIH ChestXRay dataset [wangChestXRay8HospitalScaleChest2017]; COCO (Common Objects in Context) [linMicrosoftCOCOCommon2015]; and MIDRC COVIDx CXR-4 [wuCOVIDxCXR4Expanded2023].
+22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [36]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [33]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [21], the NIH ChestXRay dataset [34]; COCO (Common Objects in Context) [18]; and MIDRC COVIDx CXR-4 [35].
 For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
   • Tabular datasets.
-After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [lucieImprovedPython2024].
+After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [11].
 For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
@@
 where $k$ is a parameter that was varied during experiments to test for effects on results.
-This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [leinsterEntropyDiversityAxiomatic2020].
-Imaging datasets were first embedded into a two-dimensional space using UMAP [mcinnesUMAPUniformManifold2020] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq.
+This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [17].
+Imaging datasets were first embedded into a two-dimensional space using UMAP [20] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq.
 7 equivalent to the root-mean-square distance (RMSD).
 The default $k$ for tabular datasets was $1$.
-3.3 Entropy calculations
-The greylock Python package [nguyen2023greylock] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1).
+3.3 Entropy and S-entropy calculations
+The sentropy Python package [23] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1).
 The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2).
-VS was calculated using the vendi‑score package [vendi-score].
+VS was calculated using the vendi‑score package [10].
 This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR.
@@
 3.4 Clustering
-HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [mcinnesHdbscanHierarchicalDensity2017] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset.
+HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [19] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset.
 HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
 3.5 Numerical experiments
-Following a previous approach [Couch:2023pav], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize.
+Following a previous approach [8], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize.
 A negative value indicates a counterexample.
@@
 4.1 LCR and VS of major imaging and tabular ML datasets
-We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits; the results below are for a random 10,000-image subset, taken for computational convenience.
+We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits.
+The results below are for a random 10,000-image subset, taken for computational convenience.
 Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
@@
 2a displays a UMAP embedding of the data, coloring each digit by its label.
-HDBSCAN detects six clusters, grouping together digits that appear visually similar (Fig.
+HDBSCAN detects six clusters, successfully grouping together digits that appear visually similar (Fig.
 2b).
-LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct (Fig.
+LCR yields the effective number of unique images after accounting for pairwise image similarity; it allows visually similar images to each contribute less to the entropy than if they were completely distinct from each other (Fig.
 2c).
 Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster.
-Taken together, these considerations suggest that LCR will approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose.
+Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose.
 Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10).
@@
 with zero pairwise similarity).
-This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease the entropy relative to a scenario in which all 10 digit classes were entirely distinct.
+This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
 In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements.
@@
 3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig.
-3b).
+3b, gray).
 Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits.
 VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues.
-The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small.
+The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig.
+3b.
 Figure 3: (a) Top 100 eigenimages for the MNIST digits dataset and (b) their eigenvalues (in the gray region), alongside the full eigenvalue spectrum. Note the log axes.
-Fig. 4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets.
+Fig.
+4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets and smaller (and sometimes no) differences for tabular datasets.
 Figure 4: LCR and VS for (a) imaging and (b) tabular datasets, sorted by VS ($q=1$; default values of $k$ for each dataset type).
@@
 […]
+The half-distance $k$ has a major impact on the correlation between LCR and VS.
 When $k$ is very small, all $z_{ij}$ approach 1, collapsing the system to a single individual for both LCR and VS (Fig.
@@
 6a and c, right).
-Only at intermediate $k$ values do LCR and VS provide independent information (Fig.
-6a and c, middle).
+Only at intermediate $k$ values do LCR and VS provide independent information (Figs.
+6a and 6c, middle).
 […]
@@
 $VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (9)
-Note that in the context of this conjecture, since we are talking about the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal.
-We prove this for $q=2,3,$ and $\infty$ in Appendix A.2.
-We consider these proofs strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not $q=2.5,4$, or $27.3$.
+Note that in the context of this conjecture, since we are considering the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal.
+We prove this conjecture for $q=2,3,$ and $\infty$ in Appendix A.2.
+We consider the existence of proofs for these several $q$ to be strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not, for example, at $q=2.5,4$, or $27.3$.
 The following lemma supports this conjecture holding for all $q$:
@@
 A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$.
-This search used a method that successfully found a counterexample to a different inequality previously [Couch:2023pav], adding further support for Conjecture 1 being true.
+The method behind this search has been successfully used to found counterexamples to other inequalities [8], adding further support for Conjecture 1 being true.
 We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
@@
 5.1 Richer measures of the information that ML datasets contain
-One way to get at this missing information is to use clustering algorithms, to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy).
-Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices.
-For example, HDBSCAN requires choosing min_cluster_size, min_samples, etc.
+One way to get at this missing information is to use clustering algorithms to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy).
+Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices about parameters; for example, HDBSCAN requires choosing min_cluster_size, min_samples, etc.
 Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
 The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters.
-One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership to that cluster and thereby contributing to the size of that cluster.
-This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster that many unique elements belong to fairly strongly is “ordinary” (Fig.
+One can view LCR as carrying fuzzy-clustering to its logical limit, with each unique element being its own cluster and the similarities giving other elements’ strength of membership in that cluster and thereby contributing to the size of the cluster.
+This is a fuzzy-clustering interpretation of the ordinariness, $Z\mathbf{p}_{i}$: a big cluster to which many unique elements belong fairly strongly would be considered “ordinary” (Fig.
 2c).
@@
 Traditional entropy accounts for frequency; LCR accounts for both frequency and similarity.
-In this way, LCR is a richer, more accurate measure of the information in a system.
-Note that, regardless of whether they are represented as effective numbers (as we [nguyen2023greylock] and others [hill1973diversity, jost2006entropy, leinsterEntropyDiversityAxiomatic2020] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [leinsterEntropyDiversityAxiomatic2020]; for example, at $q=1$ LCR never exceeds Shannon entropy.
-Similarity saps individual elements of some of their uniqueness, but in doing so, it reveals differences between datasets that traditional entropy cannot (Fig.
+In this way, LCR is a richer, more complete measure of the information in a system.
+Note that, regardless of whether they are represented as effective numbers (as we [23] and others [12, 13, 17] have advocated) or as bits of information, LCR is always bounded above by the corresponding traditional entropy [17]; for example, at $q=1$ LCR never exceeds Shannon entropy.
+Thus, similarity can be seen as sapping individual elements of some of their uniqueness.
+However, in doing so, it reveals differences between datasets that traditional entropy cannot (Fig.
 1).
@@
 1, 5c).
-Indeed, many datasets examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS.
-For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by one to two orders of magnitude, reflecting that many rows remain unclustered by hard clustering.
-In contrast, for imaging datasets, LCR and VS were typically one to two orders of magnitude lower than the number of clusters, indicating substantial similarity among clusters that hard clustering does not capture (Fig.
+Indeed, many of the datasets we examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS.
+For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute the the HDBSCAN total).
+In contrast, for imaging datasets, LCR and VS were typically 1-2 orders of magnitude lower than the number of clusters, indicating substantial similarity among elements across clusters, which hard clustering does not capture (Fig.
 2c).
-We interpret these systematic differences as evidence of the advantages of similarity‑sensitive entropy relative to hard clustering, for describing a system’s structure.
-We do not regard this as a failure of HDBSCAN; we note as a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering.
-However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search would be unlikely to meaningfully affect our findings.
+We interpret these systematic differences as evidence of the advantages of S-entropy relative to hard clustering, for describing a system’s structure.
+We do not regard this as a failure of HDBSCAN; a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering.
+However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search is unlikely to meaningfully affect the findings or conclusions presented.
 5.2 The effect of scaling the similarity measure/similarity matrix
 We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$.
-Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve upward or downward in the middle, while maintaining the anchoring at each end (at 0 and 1).
+Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve in the middle of this range, while maintaining the anchoring at each end (at 0 and 1).
 Why should one need to scale the similarity between elements in the first place?
 Should scaling be permitted at all?
-When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the similarity function is defined a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs.
-An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [aroraRepertoirescaleMeasuresAntigen2022].
-This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
-In contrast, when similarity is heuristic and not constrained by such principles, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$.
+When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the external knowledge defines and thereby constrains the similarity function a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs.
+An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [2]: this is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
+In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$.
 For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc).
-Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context and/or are maximally informative about the system.
+Therefore, in heuristic cases, scaling can serve as a reasonable dial that investigators can adjust to obtain values that make the most sense given the context/are maximally informative about the system.
 Such choices arise often in measurement.
-For example, we note that the parameter $q$ offers a loosely analogous scaling to that used in traditional (and similarity-sensitive) entropy, allowing investigators to adjust the emphasis on frequency.
-Lest choices feel too arbitrary, remember that specific integer values of $q$ correspond to well‑known statistics; for example, $q=2$ corresponds to Simpson’s index, which reflects the statistics of collision frequency.
-It is possible that alternative values of $k$ could similarly be linked to other quantities of interest.
-Investigating this possibility is left for future work.
+For example, the parameter $q$ is an analogous dial in traditional entropy (as well as S-entropy), allowing investigators to adjust the emphasis on frequency.
+As with $k$, prior information can constrain choice of $q$; for example, when the measurement has to do with collision frequency, $q=2$ should be used, consistent with $q=2$ corresponding to Simpson’s index.
+Whether specific values of $k$ might similarly be linked to known quantities of interest is left for future work.
 A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$).
@@
 Other scaling approaches may exhibit the same property.
-So: for heuristic similarity functions, which $k$ should one choose?
+For heuristic similarity functions, which $k$ should one choose?
 Qualitatively, the answer is a value roughly midway between the extremes (Fig.
@@
 Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored.
-Thus to some extent, the information in a system is in the eye of the beholder.
+Thus to some extent, from a quantitative perspective, the information in a system is in the eye of the beholder.
 6 Conclusion
@@
 First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order.
-Empirically, we observed a speed advantage of greylock over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig.
+Empirically, we observed a speed advantage of sentropy over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig.
 8).
-This performance gap may narrow or widen as codebases evolve.
-Figure 8: Compute times for all 53 datasets for LCR (using the greylock Python package [nguyen2023greylock]) vs.
+Figure 8: Compute times for all 53 datasets for LCR (using the sentropy Python package [23]) vs.
 VS (using vendi-score).
-Imaging datasets were all 10,000 images.
+Imaging datasets (blue) were all 10,000 images.
 Smaller datasets were faster.
@@
 Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality.
-The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [leinster2013magnitude]).
+The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [16]).
 Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
-And third, LCR in its various forms [chao2010phylogenetic, chaoAttributediversityApproachFunctional2019] (Table 1) is part of an extensively characterized framework [leinsterEntropyDiversityAxiomatic2020] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components.
+And third, LCR in its various forms [3, 4] (Table 1) is part of an extensively characterized framework [17] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components.
 In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class.
@@
 Measuring the full $\alpha$, $\beta$, and $\gamma$ decomposition across ML classes is a promising direction for future research.
-(The greylock implementation computes $\alpha$, $\beta$, and $\gamma$.)
-The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [nielsen2010quantum].
+(The sentropy package computes $\alpha$, $\beta$, and $\gamma$.)
+The necessary mathematics for extending this decomposition to VS exists, because VS is based on von Neumann (quantum) entropy, for which quantum mutual information, quantum relative entropy, and related quantum quantities are well understood [24].
 However, computing these quantities requires a matrix that simultaneously encodes each element and its class membership.
@@
 In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay.
-We anticipate that the deeper understanding of their relationship and properties through work such as that presented here will aid such investigations.
+We anticipate that the deeper understanding of their relationship and properties will aid such investigations.
 7 Acknowledgements

### v4

@@
-Which Similarity-Sensitive Entropy (S-entropy)?
+Which Similarity-Sensitive Entropy (Sentropy)?
 Abstract
-Shannon entropy is not the only entropy that is relevant to machine‑learning datasets, nor possibly even the most important one.
+Shannon entropy is not the only entropy that is relevant to machine-learning datasets, nor possibly even the most important one.
 Traditional entropies such as Shannon entropy capture information represented by elements’ frequencies but not the richer information encoded by their similarities and differences.
-Capturing the latter requires similarity‑sensitive entropy (S‑entropy).
-S‑entropy can be measured using either the recently developed Leinster–Cobbold–Reeve framework (LCR) or the newer Vendi score (VS).
+Capturing the latter requires similarity-sensitive entropy (“sentropy”).
+Sentropy can be measured using either the recently developed Leinster-Cobbold-Reeve framework (LCR) or the newer Vendi score (VS).
 This raises the practical question of which one to use: LCR or VS.
-Here we address this question conceptually, analytically, and experimentally, using 53 large and well-known imaging and tabular datasets.
+Here we address this question theoretically and numerically, using 53 large and well-known imaging and tabular datasets.
 We find that LCR and VS values can differ by orders of magnitude and are complementary, except in limiting cases.
-We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half‑distance” to parameterize this dependence.
-We prove that VS provides an upper bound on LCR for several values of the Rényi–Hill order parameter and present evidence that this bound holds for all values.
-We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur‑elements” or when the system that the dataset describes has a quantum‑mechanical character.
-In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, LCR is favored; nevertheless, for certain half‑distances the two methods can complement each other.
-1 Introduction
+We show that both LCR and VS results depend on how similarities are scaled, and introduce the notion of “half-distance” to parameterize this dependence.
+We prove the VS provides an upper bound on LCR for all non-negative values of the Rényi-Hill order parameter, as well as for negative values in the special case that the similarity matrix is full rank.
+We conclude that VS is preferable only when a dataset’s elements can be usefully interpreted as linear combinations of a more fundamental set of “ur-elements” or when the system that the dataset describes has a quantum-mechanical character.
+In the broader case where one simply wishes to capture the rich information encoded by elements’ similarities and differences as well as their frequencies, we propose that LCR should be favored; nevertheless, for certain half-distances the two methods can complement each other.
+I Introduction
 Entropy is the foundational quantitative descriptor of information, disorder, and uncertainty in a system and finds applications across not only science and engineering but in commerce and law [25, 44, 36, 46].
-Traditional entropy, as formulated by Shannon [30] and generalized by Rényi [28] (and later Tsallis [32]), depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig.
+Traditional entropy, as formulated by Shannon [45] and generalized by Rényi [43] and later Tsallis [47], depends exclusively on the frequency distribution of a system’s unique elements: these entropies represent the number of bits (or nats) required to encode the shape of the distribution, with skew distributions requiring fewer bits and flatter distributions requiring more (Fig.
 1).
@@
 1).
-Figure 1: The concepts of element, frequency distribution, traditional entropy (at $q=1$), similarity function, similarity measure ($z_{i}j$, with examples), similarity matrix ($Z$), and S-entropy (here, LCR, also at $q=1$).
+Figure 1: The concepts of element, frequency distribution, traditional entropy (here calculated at $q=1$), similarity function, similarity measure ($z_{ij}$, with examples), similarity matrix ($Z$), and sentropy (here calculated as LCR, also at $q=1$ for the most direct comparison).
 Entropy values are expressed in effective-number form, i.e.
-in units of effective number of images present in the dataset.
-Different similarity measures can be chosen (see Section 5.2); in this example, the similarity measure is the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
-Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or S-entropy.
+in units of the effective number of images present in the dataset.
+The choice of similarity measure is up to the investigator (see Section V.2); in this example, the similarity measure is defined as the normalized sum of the shared colors (0, 1, or 2) and shared features (outside color, inside color, and shape).
+Incorporating similarity into the calculation results in so-called similarity-sensitive entropy or sentropy.
 This was first accomplished by Leinster and Cobbold [24] and extended by Reeve and colleagues [42], in a framework we refer to as LCR [27].
@@
 similarity-insensitive) entropies because their frequency distributions are both flat, even though they may differ dramatically in composition.
-S-entropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [1, 2], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
-More recently, an alternative form of S-entropy to LCR has been described, consisting of the Vendi score [9, 26] and its variants or “cousins” [25]—hereafter collectively VS.
-This alternative raises the practical question of which form of S-entropy to choose for a given system or dataset: LCR or VS.
+Sentropy has proven to provide valuable insights into highly heterogeneous biological systems such as antibody and T-cell receptor (TCR) repertoires [4, 5], in which the vast majority of elements are unique, making the distributions heavy/long-tailed or nearly flat, such that traditional entropies are less informative.
+More recently, an alternative form of sentropy to LCR has been described, consisting of the Vendi score [14, 40] and its variants or “cousins” [39]—hereafter collectively VS.
+This alternative raises the practical question of which form of sentropy to choose for a given system or dataset: LCR or VS.
 The present work addresses this question from several angles—conceptual, empirical, and mathematical—including by measuring LCR and VS on a wide variety of medical and non-medical imaging and tabular datasets commonly used for ML research, and comparing their values.
-2 Related Work
+II Related Work
 Entropies fall into two classes: (i) traditional, i.e.
-similarity‑insensitive, and (ii) similarity‑sensitive, a.k.a.
-S-entropy.
-Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for S-entropy, by their similarity matrix as well (see Table 1).
-2.1 Traditional, similarity-insensitive entropy
+similarity-insensitive, and (ii) similarity-sensitive, a.k.a.
+sentropy.
+Although the literature often refers to “the” entropy (either as a shorthand or referring specifically to Shannon entropy), both classes are actually families of measures; individual members are distinguished by how they weight frequencies and, for sentropy, by their similarity matrix as well (see Table 1).
+II.1 Traditional, similarity-insensitive entropy
 […]
 Hill showed that exponentiating Rényi entropy yields the “effective number” of distinct elements in the dataset or system [18].
-Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ give greater weight to more frequent elements.
+Effective numbers, denoted $D_{q}$, use a parameter $q$ (identical to Rényi’s $\alpha$) such that $q=0$ counts the distinct types (by giving zero consideration to frequencies), while larger $q$ down-weight less-frequent elements.
 $D_{1}=\exp({H_{1}})$ is Shannon entropy in effective-number form.
-Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for certain $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [12].
+Hill demonstrated that several other familiar statistics correspond to special cases of $D_{q}$ for different $q$, including Simpson’s index for $q=2$ and the Berger–Parker index for $q=\infty$ [18].
 For this reason, Hill’s formulation is often described as a unifying framework.
 $D_{q}$ are known as the Hill numbers or D numbers (for “diversity”) and are given by
-$D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1.\end{cases}$ (1)
-$\mathbf{p}$ denotes the frequency distribution of unique elements.
+$D_{q}(\mathbf{p})=\exp(H_{\alpha}(\mathbf{p}))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}^{q}\right)^{\frac{1}{1-q}},&q\neq 1\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln p_{i}\right),&q=1\end{cases}$ (1)
+where $\mathbf{p}$ denotes the frequency distribution of unique elements.
 Other generalizations such as Tsallis entropy [47] also exist.
-What these traditional entropies have in common is that they are similarity‑insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
-2.2 S-entropy
-2.2.1 The Leinster-Cobbold-Reeve framework (LCR)
+What these traditional entropies have in common is that they are similarity-insensitive: they depend solely on $\mathbf{p}$ and ignore any relationships among the elements.
+II.2 Similarity-sensitive entropy: sentropy
+II.2.1 The Leinster-Cobbold-Reeve framework (LCR)
 LCR [27] extends Hill’s framework by incorporating information about the similarities and differences of the elements within the system, which traditional entropies do not.
@@
 The similarity of each element to itself is set to 1, making $Z$’s diagonal entries 1.
-The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity‑sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
-$D_{q}^{Z}(\mathbf{p};Z)=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}$ (2)
-Here $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency‑weighted average similarity of element $i$ to all the elements (including itself).
+The resulting quantities, denoted $D_{q}^{Z}$, are the exponentials of similarity-sensitive versions of the Rényi entropies $H_{\alpha}^{Z}$ and are given by
+$\begin{split}D_{q}^{Z}(\mathbf{p};Z)&=\exp(H^{Z}_{\alpha}(\mathbf{p};Z))\\ &=\begin{cases}\left(\sum_{i=1}^{n}p_{i}(Z\mathbf{p})_{i}^{\,q-1}\right)^{\!\frac{1}{1-q}},&q\neq 1,\\[6.0pt] \exp\!\left(-\sum_{i=1}^{n}p_{i}\ln(Z\mathbf{p})_{i}\right),&q=1.\end{cases}\end{split}$ (2)
+where $(Z\mathbf{p})_{i}=\sum_{j}z_{ij}p_{j}$ is the frequency-weighted average similarity of element $i$ to all the elements (including itself).
 When this average is large, element $i$ is considered ordinary, making Eq.
@@
   • attribute diversity [9] when interpreting the system as a set of attribute contributions instead of frequencies; and
-  • similarity-sensitive or similarity-aware diversity more generally [23], for which “S-entropy” is a convenient shorthand.
-LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig.
+  • similarity-sensitive or similarity-aware diversity more generally [37], for which “sentropy” is a convenient shorthand.
+LCR has proven useful for describing many complex systems whose empirical samples are uniform or close to uniform—i.e., where unique elements’ frequency distribution is flat or nearly so—a regime where traditional entropies become less informative (Fig.
 1).
-Representative applications in the life sciences include high‑throughput immunology (immunomes) [2], microbiome research (metagenomics) [23], and medical imaging [6] Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (e.g.
-images), LCR has been shown to help identify performance predictors beyond simple dataset size or class balance [6].
-2.2.2 The Vendi score and its variants (VS)
-VS constitutes a related but separate class of similarity‑sensitive entropy measures.
+Representative applications in the life sciences include high-throughput immunology (immunomes) [5], microbiome research (metagenomics) [37], and medical imaging [11].
+Specifically in ML contexts, where training sets (e.g., image collections) are often composed of unique observations (images), LCR has been shown to help identify performance predictors beyond dataset size or class balance [11].
+II.2.2 The Vendi score and its variants (VS)
+VS constitutes a related but separate class of sentropic measures.
 Like LCR, VS entropies are functions of a similarity matrix, but now the matrix has dimensions $n\times n$, where $n$ is the number of observations or total elements; for this reason we refer to it as $Z_{n}$ to distinguish it from the $Z$ used in LCR, which has one row/column per unique element.
-(When all elements are unique, $Z$ and $Z_{n}$ are the same.) We can define $Z_{n}$ in terms of $Z$ as
+(When all elements are unique, $Z=Z_{n}$.) We can define $Z_{n}$ in terms of $Z$ as
 $\left(Z_{n}\right)_{i,j}=Z_{s(i),s(j)}.$ (3)
 where $s(i)$ is the unique element of which the $i^{\text{th}}$ overall element is an instance or example.
-The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\lambda_{i}$ of $Z_{n}/n$.
-The division by $n$ normalizes $Z_{n}$ to unit trace (because as in $Z$, the diagonals of $Z_{n}$ equal 1):
-$\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\lambda_{i}\log\lambda_{i}\right)$ (4)
+The original VS is defined as the exponential of the Shannon entropy of the eigenvalues $\tilde{\lambda}_{i}=\frac{\lambda_{i}}{n}$ of $Z_{n}/n$ (where $\lambda_{i}$ are the eigenvalues of $Z$):
+$\mathrm{VS}{(\mathbf{p}_{n};Z_{n})}=\exp\!\left(-\sum_{i=1}^{n}\tilde{\lambda}_{i}\log\tilde{\lambda}_{i}\right)$ (4)
+The division by $n$ normalizes $Z_{n}$ to unit trace (because as in $Z$, the diagonals of $Z_{n}$ equal 1).
 VS is to this similarity matrix what the von Neumann entropy is to the quantum density matrix.
 Replacing the Shannon term in Eq.
-4 with a Rényi entropy of order $q$ yields the so‑called “cousins” [25] of the Vendi score:
-$\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\lambda_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
+4 with a Rényi entropy of order $q$ yields the so-called “cousins” [39] of the Vendi score:
+$\mathrm{VS}_{q}{(\mathbf{p}_{n};Z_{n})}=\left(\sum_{i=1}^{n}\tilde{\lambda}_{i}^{q}\right)^{\frac{1}{1-q}}$ (5)
 $\mathrm{VS}_{q}$ is a 1-parameter family of scores with $q$ as in $D_{q}^{Z}$.
@@
 Rényi entropy ($\alpha$)
-$H_{\alpha}(mathbf{p})=\frac{1}{1-\alpha}\log\sum_{i}p_{i}^{\alpha}$
+$H_{\alpha}(\mathbf{p})=\frac{1}{1-\alpha}\log\sum_{i}p_{i}^{\alpha}$
 Any $q$ No Info-theoretic security, Ecology, Fractal analysis, Physics
@@
 […]
-$q=1$ Yes $\lambda_{i}$ are eigenvalues of $Z_{n}=Z/N$ (positive semi-definite). ML diversity, Generative model evaluation, Ecology
+$q=1$ Yes $\tilde{\lambda}_{i}$ are eigenvalues of $Z_{n}=Z/N$ (positive semi-definite). ML diversity, Generative model evaluation, Ecology
 Cousins of the Vendi score
-$\text{VS}_{q}(\mathbf{x},\mathbf{k})=\left(\sum_{i}\lambda_{i}^{q}\right)^{\tfrac{1}{1-q}}$
+$\text{VS}_{q}(\mathbf{x},\mathbf{k})=\left(\sum_{i}\tilde{\lambda}_{i}^{q}\right)^{\tfrac{1}{1-q}}$
 Any $q$ Yes Quantum-inspired stats, Class diversity profiling
@@
 Table 1: Summary of major entropy families, their frequency-weighting forms, similarity sensitivity, and main applications.
-3 Methods
-3.1 Datasets and definitions
-The following imaging and tabular datasets were used for benchmarking LCR and VS:
+III Theorem: VS bounds LCR
+We now present our main theoretical result:
+Theorem 1.
+Let $Z$ be an $n\times n$ ($n>0$) positive semi-definite (PSD) similarity matrix, i.e.
+a PSD symmetric matrix with entries between 0 and $1$ whose diagonal entries equal $1$.
+Denote LCR and VS at order $q$ as $D_{q}(Z,p=\frac{1}{n})$ and $\text{VS}_{q}(Z)$, respectively, on the uniform distribution.
+Then for $q\geq 0$ (and for $q<0$ if $Z$ is full rank):
+$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (7)
+Following recent work [1, 17, 21, 49, 7], large language models (LLMs; especially Claude Opus 4.6) contributed to the proofs below. All results were manually checked.
+Lemma 1.
+Let $Z$ be an $n$-by-$n$ ($n>0$) non-negative symmetric matrix.
+Denote by $\lambda_{i}$ the (ordered) eigenvalues of $Z$, define $r_{i}=\sum_{j}Z_{ij}$, and let $R$ be the diagonal matrix defined by $R_{ii}=r_{i}$.
+Then $Z\preceq R$, i.e.
+RZR-Z is PSD.
+Proof.
+We must show that for any complex valued vector $v$, $v^{\dagger}\left(R-Z\right)v\geq 0$. Computing, we find
+$v^{\dagger}\left(R-Z\right)v=\sum_{i}r_{i}|v_{i}|^{2}-\sum_{i,j}v_{i}^{\dagger}Z_{ij}v_{j}$ (8)
+$=\sum_{i,j}Z_{ij}\left(|v_{i}|^{2}-v_{i}^{\dagger}v_{j}\right)$ (9)
+$=\frac{1}{2}\sum_{i,j}\left[Z_{ij}\left(|v_{i}|^{2}-v_{i}^{\dagger}v_{j}\right)+Z_{ji}\left(|v_{j}|^{2}-v_{j}^{\dagger}v_{i}\right)\right]$ (10)
+$=\frac{1}{2}\sum_{i,j}Z_{ij}|v_{i}-v_{j}|^{2}\geq 0$ (11)
+where we have used the symmetry of Z, and the final inequality follows since Z is non-negative, i.e. $Z_{ij}\geq 0$. ∎
+Lemma 2.
+Let Z be as in 1 and define $\lambda_{i}$, $r_{i}$, and $R$ w.r.t. $Z$ as in 1. Construct
+$M=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}.$ (12)
+(Note that since each $r_{i}\geq 1$, $R$ is invertible). Then $M$ is a PSD matrix with the same rank as $Z$, all of whose eigenvalues fall in the range $[0,1]$.
+Proof.
+[…]
+Note that $Z$, $R$, and $R^{-1/2}$ are real symmetric matrices, and hence Hermitian, and furthermore by assumption $Z$ is PSD.
+Consequently $M=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}$ is also PSD, as follows.
+Consider arbitrary complex vector $v$.
+Because $\left(R^{-\frac{1}{2}}\right)^{\dagger}=R^{-\frac{1}{2}}$, it is true that
+$v^{\dagger}Mv=v^{\dagger}R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}v=(R^{-\frac{1}{2}}v)^{\dagger}Z(R^{-\frac{1}{2}}v)$ (13)
+But since $Z$ is PSD, we have
+$v^{\dagger}Mv=(R^{-\frac{1}{2}}v)^{\dagger}Z(R^{-\frac{1}{2}}v)\geq 0.$ (14)
+Since $v$ was arbitrary, this proves $M$ is PSD.
+To show that the eigenvalues of $M$ are bounded above by 1, it suffices to show that $M\preceq I$, i.e.
+that IMI-M is PSD.
+This follows because $Z\preceq R$, as proved in 1, but $I-M=R^{-\frac{1}{2}}\left(R-Z\right)R^{-\frac{1}{2}}$.
+We thus find that all the eigenvalues of $M$ must be in $[0,1]$.
+An alternative demonstration that the eigenvalues are bounded above by $1$ follows from the Perron-Frobenius theorem [41, 15].
+Since $M$ is non-negative, the Perron-Frobenius theorem guarantees that $M$ has a (not necessarily unique) non-negative eigenvector, and that all non-negative eigenvectors have a single eigenvalue corresponding to the spectral radius of $M$.
+Such an non-negative eigenvector is provided by $R^{\frac{1}{2}}\mathbf{1}$ (with $\mathbf{1}$ being the all-ones vector), which we confirm by computing
+$MR^{\frac{1}{2}}\mathbf{1}=R^{-\frac{1}{2}}ZR^{-\frac{1}{2}}R^{\frac{1}{2}}\mathbf{1}=R^{-\frac{1}{2}}Z\mathbf{1}=R^{-\frac{1}{2}}r=R^{\frac{1}{2}}\mathbf{1}$ (15)
+Since the associated eigenvalue is $1$, and since we have already shown that $M$ is PSD, we find that all the eigenvalues of $M$ must be in $[0,1]$. ∎
+Lemma 3.
+Let $Z$ be as in 1. Then for real $q\geq 1$,
+[…]
+Likewise, for real $0<q<1$,
+[…]
+Proof.
+Let $M$ be as in 2.
+2 established that $M$ is PSD and has eigenvalues in $[0,1]$.
+Note that $Z=R^{\frac{1}{2}}MR^{\frac{1}{2}}$.
+By the Araki-Lieb-Thirring inequality [28, 3, 6], for any PSD matrices $A$ and $B$, and for any $p\geq 0$ and $q\geq 1$, we have:
+[…]
+Substituting $A=R^{\frac{1}{2}}$, $B=M$, and $p=1$, for $q>1$:
+[…]
+where the equality is by the cyclicity of the trace. However, because the eigenvalues of $M$ are in the range $[0,1]$ we have for $q>1$ that $M^{q}\preceq M$, meaning
+[…]
+proving the claim for $q>1$. For $0<q<1$, the direction of the Araki-Lieb-Thirring inequality reverses, yielding
+[…]
+but also $M\preceq M^{q}$ (since $x^{q}\geq x$ for $0\leq x\leq 1$ and $q\leq 1$) yielding
+[…]
+proving the claim for $0<q<1$. ∎
+Proof of 1.
+We can now prove 1 case by case.
+Case 1: $q=+\infty$
+The proof is an application of the Gershgorin circle inequality. First note that
+$\mathrm{VS}_{\infty}(Z)$ $=\frac{n}{\lambda_{\text{max}}}$ (23)
+$D_{\infty}(Z,\frac{1}{n})$ $=\frac{n}{r_{\text{max}}}$ (24)
+where $\lambda_{\text{max}}$ is the largest eigenvalue of $Z$, i.e.
+$\lambda_{\text{max}}=\displaystyle\max_{i}(\lambda_{i})$, and likewise $r_{\text{max}}=\displaystyle\max_{i}(r_{i})$.
+However, by the Gershgorin circle inequality, every eigenvalue of $Z$ must fall within a Gershgorin interval, i.e.
+in one of the intervals $Z_{ii}-\sum_{j\neq i}Z_{i,j},Z_{ii}+\sum_{j\neq i}Z_{i,j}]$.
+As such, noting that $Z_{ii}+\sum_{j\neq i}Z_{i,j}=\sum_{j}Z_{i,j}=r_{i}$, we must have that every eigenvalue $\lambda_{\mu}$ of $Z$ is bounded by $\lambda_{\mu}\leq r_{\text{max}}$.
+But from 3, this means that
+$\mathrm{VS}_{\infty}(Z)=\frac{n}{\lambda_{\text{max}}}\geq\frac{n}{r_{\text{max}}}=D_{\infty}(Z,\frac{1}{n})$ (25)
+Case 2: $q>1$ and $0<q<1$
+We now compute
+$D_{q}(Z,\frac{1}{n})$ $=\left(\sum_{i=1}^{n}\frac{1}{n}(\sum_{j}Z_{ij}\frac{1}{n})^{q-1}\right)^{\frac{1}{1-q}}$
+[…]
+But since $f(x)=x^{\frac{1}{1-q}}$ is monotonically decreasing for $q>1$ (and $\frac{1}{n^{q}}$ is a positive constant), by 3 we get
+[…]
+[…]
+On the other hand, since $f(x)=x^{\frac{1}{1-q}}$ is monotonically increasing for $0<q<1$, but the inequality from 3 reverses in this case, we get the same inequality in this case.
+Case 3: $q=1$
+We have proven the inequality for $q>1$ and $0<q<1$. Because both sides of the inequality are continuous functions of $q$ [27] [39], the result for $q=1$ follows by continuity.
+Case 4: $q=0$
+We begin by computing
+$D_{0}(Z,\frac{1}{n})=\sum_{i}\frac{1}{r_{i}}$ (28)
+[…]
+Construct $M$ as in 2. The entries of $M$ are given by
+$M_{ij}=\frac{Z_{ij}}{\sqrt{r_{i}r_{j}}},$ (30)
+with diagonal entries
+$M_{ii}=\frac{Z_{ii}}{r_{i}}=\frac{1}{r_{i}}.$ (31)
+Consequently,
+[…]
+[…]
+[…]
+proving 1 for $q=0$, completing the proof for all $q\geq 0$.
+Case 5: $q\in[-\infty,0)$ for $Z$ full rank
+[…]
+$VS_{q}(Z)\geq VS_{0}(Z)=n$ (34)
+Finally, for $q=$-$\infty$, [27] proves that the LCR diversities likewise decrease monotonically in $q$, implying
+$D_{-\infty}(Z,\frac{1}{n})\geq D_{q}(Z,\frac{1}{n})$ (35)
+Define $r_{i}=\sum_{j}Z_{ij}$. Then $r_{i}\geq 1$, since the entries of $Z$ are non-negative, and the diagonal entries (one of which contributes to $r_{i}$) are all $1$. Thus
+$D_{-\infty}(Z,\frac{1}{n})=\frac{n}{\displaystyle\min_{i}r_{i}}\leq n$ (36)
+thereby completing the proof.
+IV Experimental tests
+IV.1 Datasets and definitions
+The following imaging and tabular datasets were used for numerical experiments comparing LCR and VS:
   • Image datasets.
-22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST, Fashion‑MNIST, CIFAR‑10 and CIFAR‑100 (from the torchvision Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [36]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [33]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [21], the NIH ChestXRay dataset [34]; COCO (Common Objects in Context) [18]; and MIDRC COVIDx CXR-4 [35].
-For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset were downscaled to $250\times 250$ pixels, and images in the COCO datast were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
+22 large ($\geq 10,000$-image) medical and non-medical vision benchmarks were used: MNIST [23], Fashion-MNIST [53], CIFAR-10 and CIFAR-100 [22] (from the torchvision [31] Python package); BloodMNIST, ChestMNIST, OctMNIST, OrganAMNIST, OrganCMNIST, OrganSMNIST, PathMNIST, and TissueMNIST from the MedMNIST collection [54, 55]; the Amphibia, Insecta, Mammalia, Plantae, and Reptilia subsets from iNaturalist [48]; the computed tomography (CT) and ultrasound (US) subsets from RadImageNet [35], the NIH ChestXRay dataset [50]; COCO (Common Objects in Context) [29]; and MIDRC COVIDx CXR-4 [52].
+For computational convenience while still preserving content, images in the iNaturalist collection as well as the MIDRC COVIDx CXR-4 dataset and the NIH ChestXRay dataset were downscaled to $250\times 250$ pixels, and images in the COCO dataset were scaled to $100\times 100$ pixels; also for convenience, for each dataset, 10,000 images were selected uniformly at random for LCR and VS calculations.
   • Tabular datasets.
 After selecting for a maximum of 30 columns and 200,000 rows, the 31 most frequently downloaded datasets from the University of California Irvine Machine Learning Repository (UCIML) were retrieved [16].
-For each dataset, identifier columns and any features containing non‑numeric strings or time‑series data were removed.
-Definitions (Fig.
-1).The terms “system” and “dataset” are used interchangeably.
+For each dataset, identifier columns and any features containing non-numeric strings or time-series data were removed.
+The terms “system” and “dataset” are used interchangeably (Fig.
+1).
 The elements of image and tabular datasets are single images and single rows, respectively.
@@
 each element is also a unique element.
-3.2 Similarity matrices
+IV.2 Similarity matrices
 Similarity between each two (unique) elements $\textbf{x}_{i}$ and $\textbf{x}_{j}$ was calculated from the Euclidean distance (L2 norm) as follows:
@@
 where $k$ is a parameter that was varied during experiments to test for effects on results.
-This similarity measure is provably positive‑definite (PSD), guaranteeing valid similarity matrices for VS (see Section 6); LCR has no such restriction [17].
-Imaging datasets were first embedded into a two-dimensional space using UMAP [20] and the default $k$ taken to be $2^{-1/2}$, making the norm in Eq.
+This similarity measure was chosen for its simplicity and for being provably positive-definite (PSD), guaranteeing valid similarity matrices for VS (see Section VI); LCR has no such restriction [27].
+Imaging datasets were first embedded into a two- or three-dimensional space using uniform manifold approximation and projection (UMAP) [34] and the default $k$ taken to be $2^{-1/2}$ (varying random seeds to verify robustness), making the norm in Eq.
 37 equivalent to the root-mean-square distance (RMSD).
-The default $k$ for tabular datasets was $1$.
-3.3 Entropy and S-entropy calculations
-The sentropy Python package [23] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section 2.1).
-The package accepts a user‑provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon‑type LCR) for direct comparison with the default VS (see Section 2.2.2).
-VS was calculated using the vendi‑score package [10].
+UMAP embeddings avoid the high dimensionality of raw pixel values, which has adverse effects on distance-based algorithms [2].
+Note that UMAP projections are most often chosen to be in two dimensions.
+Empirically, UMAP embeddings generally place similar images near each other and different images further from each other (e.g.
+Figure 2), rationalizing the use of these embeddings in the literature and validating their use here.
+We present 2D UMAP results in the main text and 3D in the appendix A.
+Unlike images, the dimensionality of the tabular datasets used here is low, rationalizing direct Euclidean distance. The default $k$ for tabular datasets was $1$.
+IV.3 Entropy and sentropy calculations
+The sentropy Python package [37] was used to compute LCR as the $\gamma$ diversity; this is the effective-number form of the corresponding entropy (see Section II.1).
+The package accepts a user-provided similarity matrix and returns results for any $q$; we primarily report results for $q=1$ (Shannon-type LCR) for direct comparison with the default VS (see Section II.2.2).
+VS was calculated using the vendi-score package [14, 39].
 This implementation computes the eigenvalues of the similarity matrix, retains the positive spectrum, and evaluates the exponential of the Shannon entropy of those eigenvalues, yielding a scalar appropriate for comparison to LCR.
-Similarity matrices were calculated on a GPU-enabled workstation equipped with multi‑core CPUs.
+Similarity matrices were calculated on a GPU-enabled workstation equipped with multi-core CPUs.
 LCR and VS computations were performed on a 2024 M4 Apple Mac mini with 64GB unified memory.
-3.4 Clustering
-HDBSCAN (Hierarchical Density‑Based Spatial Clustering of Applications with Noise) [19] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset.
-HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well‑suited for heterogeneous data.
-3.5 Numerical experiments
-Following a previous approach [8], we searched for counterexamples to Conjecture 1 (Section 4.3) by numerically minimizing $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for uniform $\mathbf{p}$ using scipy.optimize.
-A negative value indicates a counterexample.
-To meet the assumptions of this conjecture, $Z$ was constructed as a gram matrix (which are always PSD), finding an $r\times n$ matrix $w$, for some integer $r$, such that $Z=w^{\intercal}w$.
-To ensure that the diagonal elements of $Z$ are equal to $1$, the rows of $w$ were $r$-dimensional unit vectors; to ensure the entries of $Z$ are non-negative, we took the entries of $w$ to be non-negative.
-We assumed a uniform distribution of unique elements.
-Thus, we parameterized $Z$ by a $r\times n$ matrix $x$, from which we defined
-$Z=w^{\intercal}w;w_{ij}=e^{x_{ij}}(\displaystyle\sum_{k}e^{2x_{kj}})^{-1/2}$ (8)
-We minimized the difference $VS_{q}(Z)-D_{q}(Z,\frac{1}{n})$ for each combination of $q\in\{-10$, 2.7-2.7, 2-2, 1-1, 0.3-0.3, 0, $0.3$, 11, $2.7$, $3.4$, 44, 55, $2\pi$, $7$, 88, $9.1$, 10}10\}, $n\in\{3,5,10,20\}$, $r\in\{2,\lfloor n/2\rfloor,n\}$ if $n>3$ and $r\in\{2,3\}$ for $n=3$ (to avoid $r=1$).
-We also varied the distribution from which the initial guess fed to scipy.optimize was drawn.
-In total, 4420 minimization were done.
-The full Python code for performing these minimizations can be found in section A.1
-3.6 Speed tests
-Run times were calculated using the Python timeit package with 5-fold replications and measured on a 2024 M4 Apple Mac mini with 64GB unified memory.
-4 Results
-4.1 LCR and VS of major imaging and tabular ML datasets
-We began by comparing LCR to VS, both at $q=1$, on a simple, well‑known imaging dataset: MNIST handwritten digits.
+IV.4 Clustering
+HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) [33] was used to obtain a non-entropic estimate of the effective number of unique elements as the number of clusters present in each dataset.
+HDBSCAN adapts to varying densities and does not require a priori specification of the cluster count, making it well-suited for heterogeneous data.
+IV.5 Speed tests
+As a practical comparison, empirical run times for sentropy and vendi-score were calculated using the Python timeit package with 5-fold replications and measured on a 2024 M4 Apple Mac Mini with 64GB unified memory.
+IV.6 LCR and VS of imaging and tabular datasets
+We first compared LCR to VS, both at $q=1$, on a simple, well-known imaging dataset: MNIST handwritten digits.
 The results below are for a random 10,000-image subset, taken for computational convenience.
-Both quantities were expressed in their effective‑number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state‑of‑the‑art clustering algorithm HDBSCAN.
+Both quantities were expressed in their effective-number forms to enable comparison with an independent benchmark: the number of clusters obtained by the state-of-the-art clustering algorithm HDBSCAN.
 Figure 2: The first 10,000 images of the MNIST digits dataset colored (a) by digit (labeled using representative images), (b) by HDBSCAN cluster, and (c) by similarity to one of the images (a “9”).
@@
 Consequently, points that lie close together count less toward the total, whereas points separated by some distance contribute more independently, even for points that happen to be in the same cluster.
-Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly because the clusters are somewhat loose.
+Taken together, these considerations suggest that LCR should approximate the number of clusters—potentially exceeding it slightly when clusters are somewhat loose.
 Indeed, the similarities among the images in this 10,000-image subset result in it having an effective number of just 12.5 images according to LCR: roughly twice the number of clusters and 25% greater than the number of classes (10).
@@
 with zero pairwise similarity).
-This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the S-entropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
+This result can be rationalized as follows: the many stylistic variations of certain digits—e.g., the different ways to write a 1, 4, or 7—increase the sentropy, while visual similarities among digits—such as the similarity between many 4s, 7s, and 9s, which cluster together per HDBSCAN—decrease it relative to a scenario in which all 10 digit classes were entirely distinct.
 In contrast, VS for this dataset was substantially higher, yielding an effective number of 95.9 elements.
-However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen‑ or “ur”‑images, each actual image in the dataset being a linear combination of these.
+However, its interpretation differs from that of LCR: the VS value indicates that the images possess the same traditional entropy as a collection of roughly 96 mutually orthogonal eigen- or “ur”-images, each actual image in the dataset being a linear combination of these.
 Fig.
-3a displays the 100 eigen‑images with the largest eigenvalues, which dominate the total entropy (Fig.
+3a displays the 100 eigen-images with the largest eigenvalues, which dominate the total entropy (Fig.
 3b, gray).
-Some of these eigen‑images resemble recognizable digits, whereas others resemble superpositions of multiple digits.
-VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen‑images at frequencies given by the eigenvalues.
-The total number of eigen‑images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig.
+Some of these eigen-images resemble recognizable digits, whereas others resemble superpositions of multiple digits that do not otherwise correspond to anything obviously familiar.
+VS is precisely the entropy of the eigenvalue spectrum, which can be interpreted as the frequency distribution of a hypothetical dataset composed of eigen-images at frequencies given by the eigenvalues.
+The total number of eigen-images equals the total number of images in the subset (10,000); the fact that the VS effective number is much smaller than this implies that only a handful of eigenvalues are large, while most are small, as confirmed by Fig.
 3b.
@@
 Fig.
-4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets and smaller (and sometimes no) differences for tabular datasets.
+4 shows LCR and VS values for 53 well known machine-learning datasets, including medical and non-medical imaging and tabular datasets, demonstrating large differences for imaging datasets (embedded in 2D UMAP projections) and smaller (and sometimes no) differences for tabular datasets (in their native space).
 Figure 4: LCR and VS for (a) imaging and (b) tabular datasets, sorted by VS ($q=1$; default values of $k$ for each dataset type).
-4.2 Trends with LCR and VS
+IV.7 Trends with LCR and VS
 Fig.
@@
 7 shows results for $q=0$ and $\infty$.
-Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging and (b) tabular datasets.
+Figure 5: Correlations among LCR, VS, and HDBSCAN for (a) imaging (2D UMAP; 10 different random seeds) and (b) tabular datasets.
 Each point represents a dataset.
@@
 Select datasets are labeled.
-(c) UMAPs for labeled imaging datasets.
+(c) UMAPs for labeled imaging datasets, colored by cluster.
 Gray points = unclustered elements.
+Figure 6: LCR vs.
+VS at a very small half-distance, an intermediate half-distance with low correlation, and large half-distance for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets.
+The scatterplots in (a) and (c) are for $q=1$.
+Red line = linear regression fit; gray line = 1:1.
+Each point represents a dataset.
+(b) and (d): vertical gray lines = the $k$ values plotted in (a) and (c).
 […]
@@
 […]
-Figure 6: LCR vs.
-VS at $q=1$ at a very small half distance, an intermediate half-distance for which the correlation is low, and a very large half distance, together with the $R^{2}$, covariance, and correlation for $q=0,1,$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets.
-For tabular datasets, the intermediate half-distance is the one simultaneously minimizing the $R^{2}$, the correlation, and maximizing the covariance at $q=1$.
-For image datasets, the intermediate half-distance does not coincide with the global minimum of the correlation at $q=1$, because VSs and LCRs already come very close to 10,000 at that global minimum, so we chose a smaller value.
-(a) and (c): red line = linear regression fit; gray line = 1:1.
-Each point represents a dataset.
-(b) and (d): vertical gray lines = $k$s for (a) and (c).
 Figure 7: LCR vs. VS for $q=0$ and $\infty$ for (a-b) imaging and (c-d) tabular datasets at their default $k$ values (11 and $\approx 2^{-1/2}$, respectively). Each point represents a dataset.
-4.3 Bounds on LCR and VS
-In Figs.
-5 and 6, VS values were consistently higher than LCR, and both measures converge to 1 as the half‑distance decreases.
-It is therefore reasonable to ask whether the inequality $VS\geq LCR$ holds universally.
-We formalized this question as the following conjecture:
-Conjecture 1.
-Let $Z$ be an $n\times n$ ($n>0$) PSD similarity matrix, i.e.
-a PSD symmetric matrix with entries between 0 and $1$ whose diagonal entries equal $1$.
-Denote LCR and VS at order $q$ as $D_{q}(Z,p=\frac{1}{n})$ and $\text{VS}_{q}(Z)$, respectively, on the uniform distribution.
-Then for all $q\in[-\infty,+\infty]$
-$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (9)
-Note that in the context of this conjecture, since we are considering the LCR diversity on a set with the uniform distribution, there is no distinction between $Z$ and $Z_{n}$ (like that made in section 2.2.2), as these are equal.
-We prove this conjecture for $q=2,3,$ and $\infty$ in Appendix A.2.
-We consider the existence of proofs for these several $q$ to be strong evidence that Conjecture 1 holds for all $q\geq 2$, since it would be surprising for VS to bound LCR at $q=2,3,$ and $\infty$ but not, for example, at $q=2.5,4$, or $27.3$.
-The following lemma supports this conjecture holding for all $q$:
-Lemma 1.
-When $Z$ is full rank but otherwise as in 1, for all $q\in[-\infty,0]$
-$VS_{q}(Z)\geq n\geq D_{q}(Z,\frac{1}{n})$ (10)
-Proof.
-[…]
-$D_{-\infty}(Z,\frac{1}{n})\geq D_{q}(Z,\frac{1}{n})$ (11)
-Define $r_{i}=\sum_{j}Z_{ij}$. Then $r_{i}\geq 1$, since the entries of $Z$ are non-negative, and the diagonal entries (one of which contributes to $r_{i}$) are all $1$. Thus
-$D_{-\infty}(Z,\frac{1}{n})=\frac{n}{\displaystyle\min r_{i}}\leq n$ (12)
-A computational search of $\approx 4,000$ examples involving similarity matrices of up to size $50\times 50$ failed to find counterexamples for any of several values of $q$, including $q=0$ and $q=1$.
-The method behind this search has been successfully used to found counterexamples to other inequalities [8], adding further support for Conjecture 1 being true.
-We present an approach to a proof for all $q$ in Appendix A.3 and present an additional proof on bounds in Appendix A.4.
-5 Discussion
+V Discussion
 Quantifying the information contained in a system is essential for virtually any complex system.
@@
 Traditional entropy is used to measure the distribution of class sizes, which often vary; this is class balance.
-However, there is no obvious way to use class balance to measure the the information contained within or between classes, which is a property of the elements themselves.
-5.1 Richer measures of the information that ML datasets contain
+However, there is no obvious way to use class balance to measure the information contained within or between classes, which is a property of the elements themselves.
+V.1 Sentropy as a measure of dataset composition
 One way to get at this missing information is to use clustering algorithms to quantify dataset substructure as the number of clusters and possibly clusters’ relative sizes (which can be measured by traditional entropy).
 Like LCR and VS, which require a choice of similarity measure and $q$, clustering algorithms require choices about parameters; for example, HDBSCAN requires choosing min_cluster_size, min_samples, etc.
-Conceptually, the main difference between clustering and S-entropy is that S-entropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
+Conceptually, the main difference between clustering and sentropy is that sentropy accounts for similarities both within and between clusters, whereas most clustering algorithms (and classifiers) lack this ability: effectively, the within-cluster (or intra-class) similarity is always 1 and the between-cluster similarity is always 0.
 The exception is fuzzy clustering, in which each element can belong to more than one cluster, with each element being parameterized by its strength of membership for one or more clusters.
@@
 1).
-How do LCR and VS compare with the number of clusters identified by HDBSCAN?
+How do LCR and VS compare with the number of clusters (as identified by HDBSCAN)?
 Predicting the relationship is difficult: large clusters whose peripheral elements are relatively distinct tend to increase the LCR/VS values, whereas strong similarities among elements—both within a cluster and across neighboring clusters—tend to reduce them (see Figs.
@@
 Indeed, many of the datasets we examined exhibited a pronounced difference in magnitude between tabular and imaging datasets when comparing the HDBSCAN cluster count with LCR and VS.
-For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute the the HDBSCAN total).
+For tabular data sets, both LCR and VS almost always surpassed the HDBSCAN cluster count, often by 1-2 orders of magnitude, reflecting that many rows remain unclustered by hard clustering (these do not contribute to the HDBSCAN total).
 In contrast, for imaging datasets, LCR and VS were typically 1-2 orders of magnitude lower than the number of clusters, indicating substantial similarity among elements across clusters, which hard clustering does not capture (Fig.
 2c).
-We interpret these systematic differences as evidence of the advantages of S-entropy relative to hard clustering, for describing a system’s structure.
-We do not regard this as a failure of HDBSCAN; a limitation of our study that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity‑sensitive entropy and clustering.
+We interpret these systematic differences as evidence of the advantages of sentropy relative to hard clustering, for describing a system’s structure.
+We do not regard this as a failure of HDBSCAN; a limitation of our study is that we did not explore a range of HDBSCAN hyperparameters, since the goal was simply to illustrate the general relationship between similarity-sensitive entropy and clustering.
 However, we note that by the “eye test,” UMAP plots validate HDBSCAN’s clustering, suggesting hyperparameter search is unlikely to meaningfully affect the findings or conclusions presented.
-5.2 The effect of scaling the similarity measure/similarity matrix
-We demonstrated that the relationship between LCR and VS depends strongly on how the similarity‑ matrix values are scaled, which we illustrated by introducing the half‑distance parameter $k$.
+V.2 The effect of scaling $Z$
+We demonstrated that the relationship between LCR and VS depends strongly on how the similarity- matrix values are scaled, which we illustrated by introducing the half-distance parameter $k$.
 Recall that similarity between two elements ranges from 0 to 1; scaling modifies this relationship by raising the values to a power, thereby shifting the curve in the middle of this range, while maintaining the anchoring at each end (at 0 and 1).
-Why should one need to scale the similarity between elements in the first place?
+Why should one need to scale the similarity between elements?
 Should scaling be permitted at all?
 When similarity is derived from external knowledge or some principle or outside requirement, we argue that scaling should not be applied: the external knowledge defines and thereby constrains the similarity function a priori, its output for any given pair is fixed, and the similarity matrix simply collects these values for all pairs.
-An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [2]: this is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
-In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean‑distance-based similarity measures used here, there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$.
+An example is binding similarity in immune repertoires, defined as the relative dissociation constants of two antibodies (or TCRs) for a given antigen (or set of antigens) [5].
+This is the unique definition that preserves biophysical additivity; arbitrary scaling would break this relationship and thereby produce a similarity matrix that is objectively incorrect.
+In contrast, when similarity is heuristic and not constrained by such knowledge, as with the Euclidean-distance-based similarity measures used here (across different embeddings), there is no a priori reason to privilege the default similarity ($k=1$) over other values of $k$.
 For instance, RMSD provides a perfectly valid similarity definition and corresponds to Euclidean distance with $k=1/\sqrt{d}$, where $d$ is the dimension of the underlying data (number of columns for tabular data, number of pixels $\times$ number of color channels for image data, 2 for the umap embedding, etc).
@@
 Such choices arise often in measurement.
-For example, the parameter $q$ is an analogous dial in traditional entropy (as well as S-entropy), allowing investigators to adjust the emphasis on frequency.
+For example, the parameter $q$ is an analogous dial in traditional entropy (as well as sentropy), allowing investigators to adjust the emphasis on frequency.
 As with $k$, prior information can constrain choice of $q$; for example, when the measurement has to do with collision frequency, $q=2$ should be used, consistent with $q=2$ corresponding to Simpson’s index.
-Whether specific values of $k$ might similarly be linked to known quantities of interest is left for future work.
-A key feature of scaling is that it interpolates similarity‑sensitive entropy between 1 (for small $k$) and traditional entropy (for large $k$).
-When $k$ is small, the off‑diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element.
-Conversely, for large $k$ the off‑diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered.
-In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective‑number form of traditional entropy equals the dataset size.
+Whether specific values of $k$ might similarly be linked to known quantities of interest is a question left for future investigation.
+A key feature of scaling is that it interpolates sentropy between 1 (for small $k$) and traditional entropy (for large $k$).
+When $k$ is small, the off-diagonal entries of the similarity matrix approach 1, implying that all unique elements are mutually similar; in the limit the system behaves as if it contained a single unique element.
+Conversely, for large $k$ the off-diagonal entries approach zero, yielding the identity matrix in the limit; the unique elements become completely dissimilar, and traditional entropy is recovered.
+In a system where the $n$ elements are all unique—typically the case for ML data sets, including the imaging and tabular data sets examined here—the effective-number form of traditional entropy equals the dataset size.
 Consequently, for such uniform systems $k$ interpolates both LCR and VS from 1 to $n$.
 Other scaling approaches may exhibit the same property.
-For heuristic similarity functions, which $k$ should one choose?
+Which $k$ should one choose?
 Qualitatively, the answer is a value roughly midway between the extremes (Fig.
 6).
-However, our quantitative findings indicate that the optimal choice is not obvious.
-We found that the correlation between LCR and VS is minimized at $k=0.2$ for tabular data sets and $3.4$ for imaging data sets—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side).
+However, our empirical studies indicate that the optimal choice is not obvious.
+We found that the correlation between LCR and VS is minimized at $k=0.2$ for the tabular data sets and $3.4$ for the imaging data sets we tested—not at any obvious definition of the midpoint of the range (defined for example as the geometric mean of the $k$ values where the correlation first falls to 0.99 on each side).
 Moreover, the shapes of these curves, with their multiple extrema, do not appear to follow a simple parametric form, making it difficult to predict precisely where the two metrics are maximally mutually informative.
 We also observed that even at the $k$ that minimizes the LCR-VS correlation, the $R^{2}$ between LCR and VS remained relatively high in absolute terms, at $0.68$.
-Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section 6).
+Perhaps this was to be expected, since both metrics quantify the effective number of elements using the same underlying similarity measure, albeit in different ways (see Section VI).
 The dependence on $k$, however, was quite strong.
-Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non‑trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale).
+Our results suggest first testing whether the entropy lies near 1 or $n$; if so, to explore, if possible, different $k$ values to delineate a non-trivial range; and finally to select a $k$ approximately in the middle of this range (on a logarithmic scale).
 Both the choice of similarity measure and the value of $k$ have effects on the resulting entropy that cannot be ignored.
 Thus to some extent, from a quantitative perspective, the information in a system is in the eye of the beholder.
-6 Conclusion
-So: LCR or VS?
-Our results support defaulting to LCR when the goal is to measure the similarity‑adjusted entropy of a system, and to use VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur‑elements/eigen‑elements, or (ii) when the system possesses a quantum‑mechanical character (see below).
-If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non‑trivial range (see Section 5.2) to maximize their independence and thereby their combined informativeness.
-We note that, quantitatively, VS will be larger than LCR for any given $k$ and $q$.
-We observed this ordering empirically and proved it for a broad class of cases, outlining a proof for the remaining situations; a computational search uncovered no counter‑examples.
-This relationship can be rationalized by recognizing that, although VS is a similarity‑sensitive entropy, its computation includes a traditional entropy component: after incorporating similarity to derive the eigen‑elements, VS is the traditional entropy of the resulting eigenvalues; traditional entropy is always greater than or equal to similarity‑sensitive entropy.
+VI Conclusion
+So, which sentropy: LCR or VS?
+Our results support defaulting to LCR when the goal is to measure the similarity-adjusted entropy of a system, and to prefer VS only in one of two special cases: (i) when the elements are usefully thought of as linear combinations of a set of mutually orthogonal ur-elements/eigen-elements, or (ii) when the system possesses a quantum-mechanical character (see below).
+If the aim is a more general multidimensional characterization, there is no clear reason to avoid employing both metrics; keeping in mind that they are likely to correlate, our findings recommend selecting $k$ within the non-trivial range (see Section V.2) to maximize their independence and thereby their combined informativeness.
+We note that, quantitatively, VS $\geq$ LCR for any given $k$ and $q$.
+We proved this relationship for all $q\geq 0$ and for $q<0$ when $Z$ is full rank.
+This relationship can be rationalized by recognizing that, although VS is a similarity-sensitive entropy, its computation includes a traditional entropy component: after incorporating similarity to derive the eigen-elements, VS is the traditional entropy of the resulting eigenvalues; traditional entropy is always greater than or equal to similarity-sensitive entropy.
 From a practical perspective, LCR has three advantages.
-First, it is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements; both calculations scale as the square of the matrix order.
-Empirically, we observed a speed advantage of sentropy over vendi-score even when all elements were unique, a situation that frequently occurs in ML datasets (Fig.
+First, its time complexity is $\mathcal{O}{(n^{2})}$ whereas computing eigenvalues—which is required to obtain the VS—has time complexity $\mathcal{O}{(n^{3})}$ [14].
+Empirically, we observed a speed advantage of the sentropy package over the vendi-score package even when all elements were unique, as is often the case in ML datasets (Fig.
 8).
+In addition, computing the LCR is less computationally expensive whenever elements appear multiple times (i.e., when $\mathbf{p}$ is not flat/uniform), because the similarity matrix for LCR has order equal to the number of unique elements, whereas VS requires a matrix whose order equals the total number of elements.
 Figure 8: Compute times for all 53 datasets for LCR (using the sentropy Python package [37]) vs.
@@
 Even symmetric similarity matrices—the most common and most natural type, satisfying $z_{ij}=z_{ji}$, so that similarity is reciprocal—are not guaranteed to be PSD, even when their entries obey the triangle inequality.
-The prevalence of such non‑PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [16]).
-Thus, PSD‑ness is irrelevant for LCR but must be checked to apply VS.
-And third, LCR in its various forms [3, 4] (Table 1) is part of an extensively characterized framework [17] that provides similarity‑sensitive analogs of mutual information, relative entropy, and cross‑entropy and offers methods for partitioning the overall similarity‑sensitive entropy ($\gamma$) into within‑class ($\alpha$) and between‑class ($\beta$) components.
-In entropic terms, $\alpha$ is the (effective‑number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class.
+The prevalence of such non-PSD cases in practice is unknown; we note that all similarity matrices examined for our datasets were PSD (a consequence of how we chose our similarity measures; this follows from Theorem 2.11 in [26]).
+Thus, PSD-ness is irrelevant for LCR but must be checked to apply VS.
+And third, LCR in its various forms [8, 9] (Table 1) is part of an extensively characterized framework [27] that provides similarity-sensitive analogs of mutual information, relative entropy, and cross-entropy and offers methods for partitioning the overall similarity-sensitive entropy ($\gamma$) into within-class ($\alpha$) and between-class ($\beta$) components.
+In entropic terms, $\alpha$ is the (effective-number form of the) joint entropy of the elements together with their class label; $\beta$ is the mutual information between the elements and the class label; and $\gamma$ is the marginal entropy of the elements irrespective of class.
 In this work we computed LCR in its $\gamma$ form; when only a single class is present, $\gamma$ coincides with $\alpha$.
@@
 In contrast, LCR requires only the similarity matrix $Z$.
-Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class‑level similarity arises purely from element‑level similarities within the classes.
+Consequently, VS needs an explicit notion of similarity both between classes and between elements, whereas in LCR any class-level similarity arises purely from element-level similarities within the classes.
 In principle, the VS formulation would permit the confusing situation in which two classes contain exactly the same individuals but nonetheless have zero similarity—a scenario that cannot occur under LCR, in which the similarity between classes is purely a function of similarities among the elements they contain.
-In sum, LCR and VS provide complementary descriptions of similarity‑sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay.
+In sum, LCR and VS provide complementary descriptions of similarity-sensitive entropy, a unifying concept that is likely to find applications in many domains in which traditional entropy is currently a mainstay.
 We anticipate that the deeper understanding of their relationship and properties will aid such investigations.
-7 Acknowledgements
+VII Acknowledgements
 This work was supported by the Gordon and Betty Moore Foundation and by the NIH under grants R01HL150394, R01HL150394-SI, R01AI148747, and R01AI148747-SI.
-Appendix A Appendix
-A.1 Python code for finding counterexamples to $VS\geq LCR$
+Appendix A Comparison to 3D UMAP embedding
+Fig.
+9 shows results of the same analysis as in the right-hand panel of Fig.
+5a but with the UMAP embeddings in three dimensions instead of two.
+This supports the lower-dimensional observations that VS and LCR tend to correlate to each other, but do contain independent information ($R^{2}=0.37$).
+Figure 9: Results from 10 replicates with different random seeds for each dataset. Legend as in Fig. 5a.
+Appendix B Other mathematical results
+1 has an interesting consequence which is not used in our proof of 1, namely
+Corollary 1.
+Let Z, $\lambda_{i}$, and $r_{i}$ be as in 1.
+Denote by $r_{(i)}$ the $i$th row-sum of $Z$ ordered from largest to smallest, so that e.g.
+$r_{(1)}=\max_{i}r_{i}$.
+Then $r_{(i)}\geq\lambda_{i}$ for all $i$ in $\{1,\ldots,n\}$
+Proof.
+This is an immediate consequence of 1 and the Weyl monotonicity theorem [51]. ∎
+This would seem a good basis for proving 1 except $VS_{q}$ depends on $\sum_{i}\lambda_{i}^{q}$, but $D_{q}\left(Z,\frac{1}{n}\right)$ depends on $\sum_{i}r_{i}^{q-1}$ rather than $\sum_{i}r_{i}^{q}$.
+It is clear, however, that 1 leads to some bound involving $VS_{q}$ and $D_{q+1}$.
+Further computation reveals that the relevant bound is
+Corollary 2.
+$\left[VS_{q}(Z)\right]^{1-q}\leq n^{\frac{1+q}{q}}[D_{q+1}]^{-q}$ (38)
+for $q>0$;
+$VS_{q}(Z)\begin{cases}\leq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}}\text{ for }0<q<1\\ \geq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}}\text{ for }q>1\end{cases}$ (39)
+Proof.
+This is a direct consequence of 1. Since $r_{(i)}\geq\lambda_{i}$, and since $x\rightarrow x^{q}$ is monotonically increasing for $q>0$, we have for each $i$ that
+$r_{(i)}^{q}\geq\lambda_{i}^{q}$ (40)
+$\Rightarrow$ (41)
+$\sum_{i}\lambda_{i}^{q}\leq\sum_{i}r_{i}^{q}$ (42)
+However, for $q\neq 1$,
+$\sum_{i}r_{i}^{q}$ $=n^{1+q}[D_{q+1}]^{-q}$ (43)
+$\sum_{i}\lambda_{i}^{q}$ $=n^{q}\left[VS_{q}(Z)\right]^{1-q}$ (44)
+thereby establishing the claim for $q\neq 1$. The $q=1$ case then follows by continuity in $q$. ∎
+Note that this bound is weaker than 3 for $q>1$, since $r_{i}^{q-1}\leq r_{i}^{q}$.
+On the other hand, for $0<q<1$ it provides an upper (rather than lower) bound on $VS_{q}(Z)$, since 1q>01-q>0, but this bound is also weaker than this simple bound that $VS_{q}(Z)<n$.
+As an illustrative example, consider the case of $q=\frac{1}{2}$.
+Then 2 leads to
+$VS_{\frac{1}{2}}\leq\frac{n^{6}}{D_{\frac{3}{2}}\left(Z,\frac{1}{n}\right)}.$ (45)
+This is clearly much weaker than $VS_{\frac{1}{2}}<n$, since $D_{\frac{3}{2}}\left(Z,\frac{1}{n}\right)\leq n$ means the right hand side is never smaller than $n^{5}$.
+In fact, this is in a sense the best case.
+Taking the inequality at general $0<q<1$, namely
+$VS_{q}(Z)\leq n^{\frac{1+q}{q(1-q)}}[D_{q+1}]^{-\frac{q}{1-q}},$ (46)
+$D_{q+1}\leq n$ means the right hand side is bounded below by $n^{\frac{1+q-q^{2}}{q(1-q)}}$.
+Minimizing the exponent (with the restriction that $0<q<1$) gives $q=\frac{1}{2}$, so this right hand side is always bounded below by $n^{5}$, where as the left hand side is bounded above by nn—a very weak bound.
+Another bound can be proven:
+Lemma 4.
+Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then
+$\mathrm{VS}_{1}(Z)\leq\frac{1}{n}\left(\sum_{i,j}Z_{ij}\right)D_{1}(Z,\frac{1}{n})$ (47)
+Proof.
+We begin by constructing the diagonal matrix $W$ according to
+$W_{ii}$ $=\frac{1}{T}\sum_{j}Z_{ij}$ (48)
+$T$ $=\sum_{i,j}Z_{ij}$ (49)
+[…]
+tr[$\tr\left[\frac{Z}{n}\left(\log\frac{Z}{n}-\log W\right)\right]\geq 0$ (50)
+$\Rightarrow$
+[…]
+However, since $W$ is diagonal we have that
+[…]
+$=-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{T}$ (53)
+$=-\sum_{i}\frac{1}{n}\left(\log\frac{\sum_{j}Z_{ij}}{n}-\log\frac{T}{n}\right)$ (54)
+$=\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (55)
+As such, we have that
+$\log\mathrm{VS}_{1}(Z)\leq\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (56)
+Exponentiating both sides, and recognizing that $e^{-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}}=D_{1}(Z,\frac{1}{n})$, we have that
+$\mathrm{VS}_{1}(Z)\leq\frac{T}{n}D_{1}(Z,\frac{1}{n})$ (57)
+proving the theorem.
+Appendix C Notes on previous versions and alternative proofs
+C.1 Proof that $VS\geq LCR$ for $q\in\{2,3\}$
+In a previous version of this paper, 1 was left as a conjecture except for $q\in{2,3,+\infty}$, and $q\in[-\infty,0]$ when $Z$ is full rank.
+The proofs for $q=+\infty$ and for $q\in[-\infty,0]$ were as currently presented in Section III.
+The previous proofs for $q\in\{2,3\}$ are presented below
+Proof of 1 in the special cases of $q=2$ and $q=3$.
+We will consider the cases separately. For convenience, in what follows, denote by $\lambda_{i}$ the eigenvalues of $Z$ and $r_{i}$ the row sums of $Z$, i.e. $r_{i}=\sum_{j}Z_{ij}$.
+At $q=2$,
+[…]
+However, since $0\leq Z_{ij}\leq 1\forall(i,j)$, it follows that $Z_{i,j}^{2}\leq Z_{ij}$ and so $\sum_{i,j}Z_{i,j}^{2}\leq\displaystyle\sum_{i,j}Z_{ij}$. Therefore,
+$\frac{n^{2}}{\sum_{i,j}Z^{2}_{i,j}}\geq\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ (59)
+However, $\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ is simply $D_{2}(Z,\frac{1}{n})$, proving the $q=2$ case.
+For $q=3$, we have that
+[…]
+$=\sum_{i,j,k}Z_{ik}Z_{kj}=\sum_{k}\left(\sum_{i}Z_{ki}\right)\left(\sum_{j}Z_{kj}\right)=\sum_{i}r_{i}^{2}.$ (61)
+This gives
+[…]
+proving the $q=3$ case. ∎
+C.2 Hoffman 1966 and a proof for integer $q\geq 2$
+A previous version of this paper conjectured the following result, noting that it would help in proving 3 (which was also stated as a conjecture).
+Theorem 2 (Hoffman 1966).
+Denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$, Let $M$ be any $n\times n$ symmetric matrix with non-negative entries, and denote by $r=M\mathbf{1}$ (i.e.
+$r_{i}=\sum_{j}M_{ij}$) the row-sums of $M$.
+Then for any integer $\alpha\geq 0$
+[…]
+Since then we have learned that this result was proved by Hoffman in [19] following up on earlier work by London in [30], which had proved the special cases of $n\leq 3$ for all integers $\alpha\geq 1$, and $5\geq\alpha\geq 1$ for arbitrary $n>0$.
+This result leads to an alternative proof of 3 for integer $q\geq 1$ (and thus a proof of 1 for integer $q\geq 2$), which is essentially what we had in mind in the earlier version of this work.
+This proof is as follows:
+Proof of 3 in the special case of integer $q\geq 1$.
+Consider that
+[…]
+since $Z$ is symmetric, $0\leq Z_{i,j}\leq 1$ for each $i,j$, and $Z^{q-1}$ is non-negative (since it is the product of q1q-1 non-negative matrices).
+But since $Z$ is symmetric and non-negative, and by assumption $q-1\geq 0$ is an integer, by 2
+[…]
+C.3 Python code to search for counterexamples
+In a previous version of this paper, we discussed a numerical search for counterexamples to 1 (partly inspred by [13]). The code used for that search is provided below.
+Note this does not search over the whole space of non-negative PSD similarity matrices.
+Rather, by construction, it only searches over completely positive matrices.
+For $n$ by $n$ matrices with $n\leq 4$, all non-negative PSD matrices are completely positive, but for $n\geq 5$ this is not the case [32].
+Given the improvements in what we have proven over that version, we no longer see a need for such a numerical counterexample search; the code is provided only for archival reasons.
 import numpy as np
@@
 evals = np.linalg.eigvalsh(Z/Z.shape[0])
-    evals = np.array([e for e in evals if e>0]) # keep only the support
+# keep only the support
+evals = np.array(
+[e for e in evals if e>0]
 if q==1:
@@
 Z = x_to_Z(x, n, r)
-    return vendi_score(Z, q) - uniform_div(Z, q)
+# Compute vendi score
+ret = vendi_score(Z, q)
+# Compute and subtract
+# uniform diversity
+ret -= uniform_div(Z, q)
+return ret
 def x_to_Z(x, n, r):
-    # map from reals to positive reals, reshape
+# map from reals to positive
+# reals, reshape
 w = (x**2).reshape(r, n)
@@
 ret = dict(min_obj)
-    ret.update({’x0’: x0, ’Z0’: x_to_Z(x0, n, r), ’Z’: x_to_Z(min_obj.x, n, r)})
+ret.update({
+’x0’: x0,
+’Z0’: x_to_Z(x0, n, r),
+’Z’: x_to_Z(min_obj.x, n, r),
 return ret
-qlist = [-10, -2.7, -2, -1, -0.3, 0, 0.3, 1, 2.7, 3.4, 4, 5, 2*np.pi, 7, 8, 9.1, 10]
+qlist = [
+-10,
+-2.7,
+-2,
+-1,
+-0.3,
+0,
+0.3,
+1,
+2.7,
+3.4,
+4,
+5,
+2*np.pi,
+7,
+8,
+9.1,
+10,
 sizes = [3, 5, 10, 20]
@@
 replicates = [i for i in range(1,6)]
-combos = [{’q’: q, ’n’: n, ’r’: r, ’rep’: rep, ’scale’: s} \
-for q in qlist for n in sizes for s in scales \
-for rep in replicates \
+combos = [{
+’q’: q,
+’n’: n,
+’r’: r,
+’rep’: rep,
+’scale’: s,}
+for q in qlist for n in sizes for s in scales
+for rep in replicates
 for r in set([x for x in [2, n//2, n] if x>1])]
@@
 pool = Pool(n_workers)
+outfile = ’counterexample_search.pkl’
 for batch in trange(n_batches):
@@
 ix_end = ix_start + n_workers
-    rows = pool.map(map_func, combos[ix_start:ix_end])
+rows = pool.map(
+map_func,
+combos[ix_start:ix_end],
 results = results + rows
 results_df = pd.DataFrame(results)
-    results_df.to_pickle(’../data/vendi_inequality_counterexample_search.pkl’)
-A.2 Proof that $VS\geq LCR$ for $q\in\{2,3,\infty\}$
-Theorem 1.
-Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then when $q\in\{2,3,\infty\}$,
-$VS_{q}(Z)\geq D_{q}(Z,\frac{1}{n})$ (13)
-Proof.
-We will consider the cases separately. For convenience, in what follows, denote by $\lambda_{i}$ the eigenvalues of $Z$ and $r_{i}$ the row sums of $Z$, i.e. $r_{i}=\sum_{j}Z_{ij}$.
-At $q=2$,
-[…]
-However, since $0\leq Z_{ij}\leq 1\forall(i,j)$, it follows that $Z_{i,j}^{2}\leq Z_{ij}$ and so $\sum_{i,j}Z_{i,j}^{2}\leq\displaystyle\sum_{i,j}Z_{ij}$. Therefore,
-$\frac{n^{2}}{\sum_{i,j}Z^{2}_{i,j}}\geq\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ (15)
-However, $\frac{n^{2}}{\sum_{i,j}Z_{i,j}}$ is simply $D_{2}(Z,\frac{1}{n})$, proving the $q=2$ case.
-For $q=3$, we have that
-[…]
-$=\sum_{i,j,k}Z_{ik}Z_{kj}=\sum_{k}\left(\sum_{i}Z_{ki}\right)\left(\sum_{j}Z_{kj}\right)=\sum_{i}r_{i}^{2}.$ (17)
-This gives
-[…]
-proving the $q=3$ case.
-For $q=\infty$, we invoke the Gershgorin circle inequality. We first note that
-$\mathrm{VS}_{\infty}(Z)$ $=$ $\frac{n}{\lambda_{\text{max}}}$ (19)
-$D_{\infty}(Z,\frac{1}{n})$ $=$ $\frac{n}{r_{\text{max}}}$ (20)
-where $\lambda_{\text{max}}$ is the largest eigenvalue of $Z$, i.e.
-$\lambda_{\text{max}}=\displaystyle\max_{i}\lambda_{i}$, and likewise $r_{\text{max}}=\displaystyle\max_{i}r_{i}$.
-However, by the Gershgorin circle inequality, every eigenvalue of $Z$ must fall within a Gershgorin interval, i.e.
-in one of the intervals $Z_{ii}-\sum_{j\neq i}Z_{i,j},Z_{ii}+\sum_{j\neq i}Z_{i,j}]$.
-As such, noting that $Z_{ii}+\sum_{j\neq i}Z_{i,j}=\sum_{j}Z_{i,j}=r_{i}$, we must have that every eigenvalue $\lambda_{\mu}$ of $Z$ is bounded by $\lambda_{\mu}\leq r_{\text{max}}$.
-But this means that
-$\mathrm{VS}_{\infty}(Z)=\frac{n}{\lambda_{\text{max}}}\geq\frac{n}{r_{\text{max}}}=D_{\infty}(Z,\frac{1}{n})$ (21)
-A.3 Approach to a proof of $VS\geq LCR$ for all $q$
-A proof for 1 at all $q$ except for $1$ (which could perhaps be filled in using continuity considerations) would follow from the following:
-Conjecture 2.
-Let $Z$ be as in 1, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for $q>1$,
-[…]
-Likewise, for $q<1$,
-[…]
-At $q>1$, 2 would in turn be proven by the following conjecture:
-Conjecture 3.
-Let $M$ be any $n\times n$ symmetric matrix whose entries obey $0\leq M_{ij}\leq 1$, and denote by $\mathbf{1}$ the $n$ dimensional vector each of whose entries are $1$. Then for any $\alpha\geq 0$
-$\mathbf{1}^{\intercal}M^{\alpha}\mathbf{1}\leq\mathbf{1}^{\intercal}\left(M\mathbf{1}\right)^{\alpha}$ (24)
-A.4 Other bounds
-Beyond the conjectures upper bound of uniform diversity-with-similarity by Vendi score, we are able to prove the following bound in the opposite direction:
-Theorem 2.
-Let $Z$, $\text{VS}_{q}(Z)$, and $D_{q}(Z,\frac{1}{n})$ be as in 1. Then
-$\mathrm{VS}_{1}(Z)\leq\frac{1}{n}\left(\sum_{i,j}Z_{ij}\right)D_{1}(Z,\frac{1}{n})$ (25)
-Proof.
-We begin by constructing the diagonal matrix $W$ according to
-$W_{ii}$ $=$ $\frac{1}{T}\sum_{j}Z_{ij}$ (26)
-$T$ $=$ $\sum_{i,j}Z_{ij}$ (27)
-[…]
-tr[$\tr\left[\frac{Z}{n}\left(\log\frac{Z}{n}-\log W\right)\right]\geq 0$ (28)
-$\Rightarrow$
-[…]
-However, since $W$ is diagonal we have that
-[…]
-$=$ $-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{T}$ (31)
-$=$ $-\sum_{i}\frac{1}{n}\left(\log\frac{\sum_{j}Z_{ij}}{n}-\log\frac{T}{n}\right)$ (32)
-$=$ $\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (33)
-As such, we have that
-$\log\mathrm{VS}_{1}(Z)\leq\left(-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}\right)+\log\frac{T}{n}$ (34)
-Exponentiating both sides, and recognizing that $e^{-\sum_{i}\frac{1}{n}\log\frac{\sum_{j}Z_{ij}}{n}}=D_{1}(Z,\frac{1}{n})$, we have that
-$\mathrm{VS}_{1}(Z)\leq\frac{T}{n}D_{1}(Z,\frac{1}{n})$ (35)
-and so the theorem is proved.
+results_df.to_pickle(outfile)